### v1

# EquiContact: A Hierarchical SE(3) Vision-to-Force Equivariant Policy for Spatially Generalizable Contact-rich Tasks

Abstract
This paper presents a framework for learning vision-based robotic policies for contact-rich manipulation tasks that generalize spatially across task configurations. We focus on achieving robust spatial generalization of the policy for the peg-in-hole (PiH) task trained from a small number of demonstrations. We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant ACT, G-CompACT). G-CompACT operates using only localized observations (geometrically consistent error vectors (GCEV), force-torque readings, and wrist-mounted RGB images) and produces actions defined in the end-effector frame. Through these design choices, we show that the entire EquiContact pipeline is $SE(3)$-equivariant, from perception to force control. We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance. Real-world experiments on PiH tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles. The experimental videos can be found on the project website: https://sites.google.com/berkeley.edu/equicontact
I Introduction
Imitation learning has recently shown significant success in expanding the capabilities of machine learning in real-world robotics applications [1, 2]. In the early stages of robot learning, many methods formulated manipulation as sequences of keyframe-based pick-and-place actions [3, 4]. More works have started to produce a continuous set of actions directly from vision inputs [5, 6, 7]. Similar to the trend seen in large language models (LLMs), there is a growing belief that large-scale data can unlock generalizable, vision-based policies for robotics [8]. This has led to massive efforts to build large datasets [1] for training policies with general knowledge.
However, such policies often lack spatial generalizability and therefore require a large amount of data to learn robust policies. As described in [9], both action chunking transformers (ACT) [6] and diffusion policy (DP) [5] exhibit near-linear performance growth as the demonstration dataset size increases, suggesting that the trained policies do not inherently generalize well to new spatial configurations.
An alternative line of recent research focuses on leveraging symmetry—particularly equivariance—to enhance spatial generalizability, thereby improving sample efficiency during training [10, 11]. This approach requires fewer data but comes with its own challenges. Equivariant neural networks, being of a more specialized nature, are often not as well-developed and are more computationally intensive than their non-equivariant counterparts, making real-time and large-scale deployment more difficult. As a result, it becomes more attractive for users to use standard models trained with massive datasets in many instances.
[…]
In this paper, we propose EquiContact, a hierarchical SE(3) vision-to-force equivariant policy for spatially generalizable contact-rich tasks. Our proposed method consists of two main components. The first is a high-level planner consisting of a Diff-EDF model [14], which uses global point cloud data to provide a local reference frame for placing the peg relative to the hole. The second, low-level compliant visuomotor policy is a variant of ACT [6], which we refer to as Geometric Compliant ACT (G-CompACT). This policy handles detailed motion and contact interaction using force-torque feedback and RGB inputs from wrist-mounted cameras. A key design feature of G-CompACT is that it only relies on local information—specifically, the force-torque signal in the end-effector frame, a geometrically consistent error vector (GCEV) [11], and wrist camera inputs. The output of G-CompACT is the desired pose and admittance gains, which are then sent to the geometric admittance controller (GAC) module to execute compliant control. We emphasize that our contribution is on the structural framework rather than the specific choice of algorithms. For example, one could replace the Diff-EDF with other equivariant methods, such as ET-SEED [15], or replace the ACT with the DP or their variants, e.g., Diffusion Transformer (DiT) [16]. The main contributions of this paper are as follows:
  1. 1.
    We propose EquiContact, a hierarchical, provably $SE(3)$-equivariant policy from point clouds and RGB inputs to interaction forces for executing contact-rich tasks.
  2. 2.
    We demonstrate that EquiContact achieves near-perfect success rates and spatial generalizability in real robot experiments involving tight insertions.
  3. 3.
    We identify three key principles for spatially generalizable contact-rich manipulation: Compliance, Localized Policy (Left invariance), and Induced Equivariance. These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [10].
From these key principles, we propose a general framework to enhance the spatial generalization and interpretability of vision-based policies, namely, “anchoring localized policy on globally estimated reference frame.” We emphasize that our work provides complementary insights to recent trends of robot learning [1, 8, 16, 2] that aim to build generalist policies from a large-scale demo dataset. Our principles offer structural guidelines to improve spatial generalizability through $SE(3)$ equivariance.
II Related Works
Visuomotor Servoing Methods Recently, generative modeling has become mainstream in realizing visuomotor servoing policies. Particularly, there are two dominant methods for visuomotor servoing: Action Chunking with Transformers (ACT) [6] and Diffusion Policy (DP) [5]. ACT utilizes the conditional variational autoencoder (CVAE) method as a generative model, whereas DP employs denoising diffusion as its generative model. ACT and DP have been extended to other approaches, including compliance and force-reactive behaviors [17, 18, 19], as well as structural improvements [20, 16, 21]. Our work is most closely related to CompliantACT (CompACT) [17], which integrates compliant control for visuomotor policies. We have significantly improved CompACT by incorporating a provable $SE(3)$ equivariant structure.
Equivariant Methods Equivariant methods aim to leverage the inherent symmetries of robot tasks, typically modeled as group transformations such as $SE(3)$, to improve the sample efficiency of imitation learning. Early approaches, such as Neural Descriptor Fields (NDFs) [22] and Transporter Networks [3], incorporated equivariance to $SE(3)$ and $SO(2)$ transformations, utilizing scene point clouds and top-down views, respectively. Equivariant Descriptor Fields (EDFs) [23] reformulated the NDF approach into a fully end-to-end learning method with $SE(3)$ bi-equivariance property, which is further improved by Diffusion-EDF (Diff-EDF) [14] to improve training and inference performance. Other equivariant models that extract desired keyframes from point clouds include [24]. 3D Equivariant extensions of DP [25, 26, 15] and Flow-Matching [27, 28] policies have been explored in recent literature and combine $SE(3)$ equivariant observation encoders with appropriately modified diffusion/flow-matching processes. Compared to recent works of [29, 30], we generalize the table-top manipulation to contact-rich tasks while achieving full $SE(3)$ equivariance from vision to force/control level, bridging vision, control, and force interaction under a unified framework.
Figure 1: We propose an EquiContact, a hierarchical, provably $SE(3)$ vision-to-force equivariant policy for spatially generalizable contact-rich tasks. The proposed EquiContact consists of G-CompACT and Diffusion-EDF. The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen scenarios during evaluation.
III Problem Definition
In this paper, we aim to identify the key structural components of a policy capable of executing spatially generalizable, contact-rich manipulation tasks. As mentioned earlier, our proposed building blocks to achieve $SE(3)$ vision-to-force equivariant policy are: 1. Compliant Control, 2. Localized Policy, and 3. Induced Equivariance - which we will experimentally validate through specific data collection and evaluation setups, as detailed below.
III-A Problem Setup
Our primary focus is to provide a learning-based solution for a peg-in-hole task,a classic representative of contact-rich, force-based robotic manipulation and assembly tasks, using vision. Unlike prior work such as [11], which assumes that the desired hole pose is known and that the robot starts with a pre-grasped peg, we consider a more general scenario. The robot first needs to pick up the peg and then insert it into the hole, as shown in Fig. 2, relying solely on vision and proprioception. We assume only that the peg is upright in the scene. Following the reasoning that every manipulation task can be regarded as a sequence of pick-and-place, we formulated peg-in-hole insertion as a compliant pick-and-place task. However, as highlighted in the earlier discussion, naively using the pick-and-place approach is insufficient because the high-level vision planner often lacks the accuracy required for precise and contact-rich tasks, which is $1\mathrm{mm}$ in our case (see Fig. 2).
To accommodate this, we propose a low-level policy that:
  • provides real-time visual feedback to correct errors from the coarse high-level planner,
  • handles fine force-based interactions through compliance,
  • and achieves provable $SE(3)$-equivariance.
Assuming that a high-level vision planner that returns approximate desired end-effector pick-and-place poses is readily available, we focus on developing equivariant compliant placing, i.e., an insertion policy via imitation learning, and regard picking as a spatial case of placing.
To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location. This enables the policy to perform nearly perfectly in the trained scenario. We then evaluate benchmark and proposed methods—trained solely on these limited demonstrations—across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented. $2$ external cameras with calibrated extrinsics and $2$ wrists cameras are installed. (Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance. (Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
IV Solution Approach
We now describe our proposed EquiContact policy, which integrates a high-level vision-based planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant control Action Chunking Transformer (G-CompACT). G-CompACT itself consists of a behavior cloning module built on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC). As discussed in Section III, because of space constraints, we focus on the insertion task in this paper, which will later be extended to include the picking task to complete the full pipeline implementation. We note that the Diff-EDF and G-CompACT are trained separately but are combined as a whole pipeline during inference time, as shown in Fig. 1.
IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
Diff-EDF [14] is a diffusion-based manipulation policy with a bi-equivariant structure on $SE(3)$ transformations for pick-and-place tasks. Diff-EDF takes two point clouds as input: the scene point cloud $\mathcal{O}^{scene}$ captured by two external RGBD cameras with calibrated extrinsics (Fig. 2), and the gripper point cloud $\mathcal{O}^{grasp}$ in the end-effector frame, with or without a grasped object.
Diff-EDF outputs a pose $g_{{}_{EDF}}\in SE(3)$ representing the estimated target pose of the object of interest, which in our case is the hole position and is defined as:
$g_{{}_{EDF}}=f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$ (1)
where the neural network $f_{\theta_{1}}$ implements the mapping. Due to the left-equivariant property, it follows that:
$g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
Let $\mathcal{O}^{ref}\subset\mathcal{O}^{scene}$ denote the subset of points corresponding to the object of interest (e.g., the hole structure, see also Fig. 3). The model is designed to satisfy the following left-equivariance property on the point cloud characterization of the target object of interest:
Assumption 1 (Left-Equivariance of Diff-EDF).
$\begin{split}g_{{}_{EDF}}&=f_{\theta_{1}}(\mathcal{O}^{ref},\mathcal{O}^{grasp})\\ &\implies g_{l}\cdot g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref},\mathcal{O}^{grasp}),\end{split}$ (2)
To meet this assumption, we randomize the pose of the hole-platform assembly during training, both translationally and rotationally, with visual distractors in the background. This encourages the model to focus on $\mathcal{O}^{ref}$, not the rest of the objects in the scene.
While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and have high robustness to left $SE(3)$ transformations.
To train Diff-EDF, the scene and grasp point clouds are collected together with the target reference frames, which represent the desired poses of the end-effector for pick-and-place operations. $15$ demonstrations were collected for the Diff-EDF: $12$ samples of the flat platform and $3$ samples of the tilted platform, both translationally and rotationally randomized. The training process of Diff-EDF follows the procedure in [14].
While the original Diff-EDF supports right-equivariance to handle transformations between the gripper and the grasped object, we did not utilize this feature. In practice, we observed that arbitrary peg transformations relative to the gripper introduced additional challenges. First, grasping the peg based on an arbitrary transformation leads to imprecise grasps, resulting in object slippage during contact. Second, the gripper-object rotation must be estimated continuously in real-time for G-CompACT to work under slippage. Instead, we enforced a consistent grasp orientation and relied only on the left-equivariance property. We refer the readers to [14, 10] for full details of Diff-EDF and to Supplementary Material  SI for implementation and heuristics. We emphasize again that Diff-EDF can be replaced by any point cloud-based method that outputs an approximate goal pose for (compliant) placing.
IV-B Geometric Compliant control Action Chunking with Transformers (G-CompACT)
IV-B1 Geometric Admittance Control (GAC)
We implement the geometric impedance control (GIC) proposed in [12, 13] in the geometric admittance control (GAC) setup [11]. We first note that we follow the notation used in [11]. Let the end-effector pose be denoted as $g\in SE(3)$ in a homogeneous matrix representation, or simply $g=(p,R)$, where $p\in\mathbb{R}^{3}$ is a position of the end-effector and $R\in SO(3)$ is a rotation matrix of the end-effector. Given the desired end-effector pose $g_{d}=(p_{d},R_{d})$, the desired end-effector dynamics for the GAC setup is given as follows:
$M\dot{V}^{b}+K_{d}V^{b}+f_{{}_{G}}=F_{e},$ (3)
where $M\in\mathbb{R}^{6\times 6}$ is symmetric positive definite desired inertia matrix, $K_{d}\in\mathbb{R}^{6\times 6}$ symmetric positive definite damping matrix, $F_{e}\in\mathbb{R}^{6}$ is external wrench applied to the end-effector in end-effector body frame and $V^{b}\in\mathbb{R}^{6}$ is a body-frame end-effector velocity. Further, $f_{{}_{G}}=f_{{}_{G}}(g,g_{d},K_{p},K_{R})\in\mathbb{R}^{6}$ is a elastic wrench given by:
$\begin{split}f_{{}_{G}}=\begin{bmatrix}f_{p}\\ f_{R}\end{bmatrix}=\begin{bmatrix}R^{T}R_{d}K_{p}R_{d}^{T}(p-p_{d})\\ (K_{R}R_{d}^{T}R-R^{T}R_{d}K_{R})^{\vee}\end{bmatrix},\end{split}$ (4)
where $\hat{f}_{{}_{G}}\in se^{*}(3)$, with $K_{p},K_{R}\in\mathbb{R}^{3\times 3}$ symmetric positive stiffness matrices for the translational and rotational dynamics, respectively, and $(\cdot)^{\vee}$ denotes the vee-map. By a slight abuse of notation we will frequently denote as $f_{{}_{G}}\in se^{*}(3)$ to emphasize that it is an elastic wrench on $SE(3)$, since $f_{{}_{G}}$ is isomorphic to $\hat{f}_{{}_{G}}$ [13].
Given the desired dynamics (3), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (5)
where $T_{s}$ is a sampling time and $\hat{(\cdot)}$ denotes a hat-map. For the details of the GIC/GAC, we refer to [12, 11]. The admittance control loop is implemented at a $200\mathrm{Hz}$ rate using ROS2. We refer to Supplementary Material SI for detailed implementation tips.
IV-B2 G-CompACT
Action Chunking with Transformer (ACT) is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks, originally introduced in [6]. While the original ACT formulation encodes robot proprioception using joint positions, this representation lacks spatial equivariance, limiting generalization. Instead, we adopt a proprioceptive encoding based on the end-effector pose, following the approach in [17].
In our previous work [11], we proposed a recipe for spatially equivariant policies, which involves a left invariance property and policy representation in the end-effector body frame. Following this principle and recent developments on relative action representations (similar to [29]), we structure our G-CompACT policy to take observations composed of: (1) Geometrically Consistent Error Vector (GCEV) proposed in [11], (2) FT sensor in the end-effector frame to capture contact behaviors, and (3) RGB images from wrist cameras. The model returns as action sequences: (1) relative pose from the current end-effector frame, and (2) admittance gains.
The GCEV $e_{{}_{G}}(g,g_{{}_{EDF}})$ is defined as
$e_{{}_{G}}(g,g_{{}_{EDF}})=\begin{bmatrix}R^{T}(p-p_{{}_{EDF}})\\ (R_{{}_{EDF}}^{T}R-R^{T}R_{{}_{EDF}})^{\vee}\end{bmatrix},$ (6)
where $g=(p,R)\in SE(3)$ is a current end-effector pose and $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF. The derivation of the GCEV $e_{{}_{G}}$ is presented in [12]. We note that the GCEV denotes an error vector on the end-effector body frame as emphasized in [11].
The model outputs a chunk of $N$ relative pose actions:
$\bar{g}_{d}(k+i)=g(k)\cdot g_{rel}(k+i),\quad i=1,\cdots,N,$ (7)
where each $g_{rel}(k+i)\in SE(3)$ is the $i$th relative pose from the ACT model and $\bar{g}_{d}(k+i)$ is the $i$th temporary desired pose signal. A temporal ensembling mechanism is then applied to $\bar{g}_{d}(k+i)$ to compute the desired pose $g_{d}(k)$ for GAC. To respect the manifold structure of $SE(3)$, we use the scipy implementation for weighted averaging on the rotation matrix to get $R_{d}(k)$ and standard weighted averaging on positions for $p_{d}(k)$ separately.
Our proposed G-CompACT also outputs the admittance gains matrices $K_{p}$ and $K_{R}$ in the GAC elastic wrench in Eq. (4). Unlike CompACT, the gain matrices are represented in the desired frame of the end-effector [13], removing the necessity of utilizing full-gain matrix representations and coordinate transformations to deal with spatial equivariance. Therefore, a diagonal matrix representation in $K_{p}$ and $K_{R}$ is sufficient to conduct the PiH task successfully, without requiring a Cholesky factorization to represent the gains, as in CompACT. Since the diagonal component representation of matrices is equivalent to a vector, the standard temporal ensemble can be applied. Similar to the relative pose action, the ACT provides temporary gain signals $\bar{K}_{p}$ and $\bar{K}_{R}$, followed by a temporal ensemble to obtain $K_{p}$ and $K_{R}$.
Formally, let the G-CompACT method be $\pi_{\theta}$ such that
$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq\{(g_{rel}(k+i),\bar{K}_{p}(k+i),\bar{K}_{R}(k+i))\}_{i=1}^{N},\\ o(k)&\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (8)
where $\bar{K}_{p}$ and $\bar{K}_{R}$ denote the temporary admittance gains, $a(k)$ denotes the actions and $o(k)$ denotes the observation at time step $k$, $e_{{}_{G}}$ is GCEV defined in (6), $F_{e}$ is force-torque sensor reading on the end-effector frame, and $I_{w}=\{I_{w,1},I_{w,2}\}$ where $I_{w,1}$ and $I_{w,2}$ denote the images from the wrist cameras #1 and #2, respectively (Fig. 2).
Since the images $I_{w}$ are fed to the transformer decoder after being passed to the visual encoder structure, one can further consider the G-CompACT as
$\begin{split}a(k)&=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)(k),\quad\text{where}\;z(k)=\mu_{\phi}(I_{w}).\end{split}$ (9)
where $\mu_{\phi}$ denotes the visual encoder, e.g., ResNet, and $z$ is the latent variable from the visual encoder.
We now introduce the following assumption regarding the left-invariant visual representation.
Assumption 2 (Left-invariant Visual Features).
The features from visual encoder $\mu_{\phi}$ are left invariant, i.e.,
$z(k)=\mu_{\phi}(g_{l}\circ I_{w})=\mu_{\phi}(I_{w}),\;\;\forall g_{l}\in SE(3).$ (10)
Note that we refer to $g_{l}$ as a left-group action [10]. The left-group action applied on the wrist-camera images is illustrated in Fig. 3. The meaning of the visual representation $z$ being left-invariant is that the vision encoder $\mu_{\phi}$ is trained to focus only on group action invariant features, such as the flat surface surrounding the hole on the platform. The satisfaction of Assumption 2 can be challenging. To satisfy this assumption, we have designed our platform and the surface surrounding the hole assembly so that there are sufficient surface features and the cameras primarily see the surface surrounding the hole, not the lower parts of the platform. This engineering choice was somewhat ad hoc, as there is no guarantee or sufficient inductive bias to encourage the desired behavior. We will demonstrate in later experimental results that this assumption may not apply in certain cases.
Figure 3: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$. As the left group action is applied on the end-effector, the wrist cameras start to see not only the optical table but also the outer backgrounds. $\{s\}$ denotes a spatial frame.
The following proposition shows the left-invariance of the G-CompACT method in the end-effector frame.
Proposition 1 (Left-invariance of G-CompACT).
Suppose that the Assumption 2 holds. Then,
$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi_{\theta_{2}}(o(k)).$ (11)
Proof.
The left-translated observation signals $o(k)$ reads that:
$g_{l}\circ o(k)=(g_{l}\circ e_{{}_{G}},g_{l}\circ F_{e},g_{l}\circ I_{w}).$ (12)
As was shown in Lemma 1 of [11], the GCEV $e_{{}_{G}}$ is left invariant as
$g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}})=e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}})=e_{{}_{G}}(g,g_{{}_{EDF}}).$
The force-torque sensor values are left-invariant because they are already defined with respect to the end-effector frame [11], and the visual representation vectors satisfy left invariance due to Assumption 1. Combining all these properties, it follows that
$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)=\pi_{\theta_{2}}(o(k)),$ (13)
which shows the left invariance of the G-CompACT policy on the end-effector frame. ∎
In what follows, we demonstrate that the pose signal produced by G-CompACT is equivariant when described in the spatial frame.
Corollary 1 ($SE(3)$ (Left) Equivariance of G-CompACT).
Let the output of G-CompACT in the spatial frame be $\bar{g}_{d}$ via (without considering the outputs related to gains)
$\bar{g}_{d}=g\!\cdot\!\pi_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (14)
Then,
$g_{l}\bar{g}_{d}$ $=g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w})$
$=g_{l}g\cdot\pi_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (15)
Since the temporary desired pose on the spatial frame signal, $\bar{g}_{d}$, is $SE(3)$ equivariant, the temporal aggregated signal $g_{d}$ is also equivariant. Therefore, by combining the G-CompACT with a spatial frame representation and temporal ensemble, we will now define the final form of G-CompACT as $\hat{\pi}_{\theta_{2}}$, which is given by:
$(g_{d},K_{p},K_{R})=\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (16)
and satisfies the following equivariance property:
$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\hat{\pi}_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (17)
To train a G-CompACT, we collect the expert demonstration using teleoperation on the fixed pose of the platform. During data collection, the expert teleoperator monitors the task’s progress, makes real-time movement commands via a SpaceMouse, and adjusts the admittance gains using keyboard input to switch between the following predefined gain modes: low-gain mode, high-gain mode, insertion mode, and contact mode. The low/high gain mode has low/high gains in all directions, the insertion mode has high gains in the $z$ direction of the end-effector frame and low gains elsewhere. Finally, the contact mode has low gains in the $z$ direction and high gains elsewhere. We collected $86$ demonstrations to train a near-perfect policy.
Since we know the fixed pose of the platform a priori, e.g., a ground-truth reference frame, the GCEV vector can be calculated for the training process. Nevertheless, the reference frame needs to be estimated via Diff-EDF (as $g_{{}_{EDF}}$) during the inference stage, which may have non-negligible errors. To handle this issue, we have added noise to the reference frame $g_{ref}$ to calculate $e_{{}_{G}}$ during dataset preprocessing. This provides the model with an inductive bias to primarily rely on $e_{{}_{G}}$ values for rough alignment and rely on vision feedback for fine-grained motion. The rest of the training follows the standard imitation learning pipeline.
IV-C EquiContact
The proposed EquiContact method comprises the high-level Diff-EDF, which serves as a high-level vision planner that provides the reference frame to be fed to GCEV, and the low-level G-CompACT, which handles fine-grained movement and contact interaction during insertion using real-time vision and force feedback. The overall pipeline of the EquiContact is presented in Fig. 1. Here, we demonstrate that the proposed EquiContact possesses the equivariance property. Let EquiContact be written as $h_{\theta}$ so that $h_{\theta}(g,g_{ref},F_{e})\mapsto f_{{}_{G}}$, i.e., $h_{\theta}:SE(3)\times SE(3)\times se^{*}(3)\to se^{*}(3)$.
Proposition 2.
The EquiContact Policy $h_{\theta}$ is equivariant if it is described relative to the spatial frame.
Proof.
Suppose that the Assumption 1 and 2 hold. Let the object of interest, e.g., a peg for the picking task and a hole for the placing task, be observed by $\mathcal{O}^{ref}$, $I_{w}$ with its pose given by $g_{ref}$, so that the left-translated $g_{l}\cdot g_{ref}$ is observed by $g_{l}\circ\mathcal{O}^{ref}$ from the point cloud, and $g_{l}\circ I_{w}$ by the left-translated end-effector attached wrist camera as described in Fig. 3.
First, notice that $h_{\theta}$ can be fully written as
$h_{\theta}(g,g_{ref},F_{e})$ $=f_{{}_{G}}(g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,f_{\theta_{1}}(\mathcal{O}^{ref})),F_{e},I_{w}))$ (18)
$=f_{{}_{G}}(g,\underset{=(g_{d},K_{p},K_{R})}{\underbrace{\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w})}}).$
Then, when both $g$ and $g_{ref}$ undergoes a left transformation $g_{l}$, from Assumption 1 and Corollary 1, the following holds:
$h_{\theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})$
$=f_{{}_{G}}(g_{l}g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g_{l}g,f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref})),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
$=f_{{}_{G}}(g_{l}g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
$=f_{{}_{G}}(g_{l}g,g_{l}g_{d},K_{p},K_{R})=f_{{}_{G}}(g,g_{d},K_{p},K_{R})$
$=h_{\theta}(g,g_{ref},F_{e}).$ (19)
We note that the second-last equation ($SE(3)$ left-invariance of the elastic wrench) comes from Lemma 1 of [11]. Finally, from the result of Proposition 2 of [11], it follows that
$h^{s}_{\theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})=\text{Ad}_{g_{l}^{\text{-}1}}^{T}h_{\theta}^{s}(g,g_{ref},F_{e}),$ (20)
where Ad is an Adjoint operator, $\text{Ad}_{g_{l}^{\text{-}1}}^{T}$ is a representation for $se^{*}(3)$ (wrench) domain for the group action $g_{l}$ [11], and superscript $s$ denotes a vector represented in the spatial frame. ∎
IV-D Extensions to Pick Tasks
So far, we have described our method in terms of the insertion (placement) task. The proposed method can be extended to pick tasks in the same manner. The Diff-EDF can be utilized to obtain the pick reference frame, which is used for $e_{{}_{G}}$ for the picking G-CompACT. The picking G-CompACT is trained in such a way that the manipulator grasps a peg in a fixed, aligned pose, which helps EquiContact bypass the right-equivariance issue. For G-CompACT, the FT sensor values are not utilized as one of its observations, and it does not output the admittance gains; instead, it uses fixed gains.
V Experimental Results and Discussions
V-A Benchmark Results
To validate the effectiveness of our proposed EquiContact framework and its underlying design principles, we compare EquiContact against three benchmark approaches: ACT with world-frame observation and action representation, executed with and without GAC, and CompACT [17]. We note that the GAC serves as a standard admittance control when the platform’s pose is fixed, i.e., untilted. The observation and action representation of the benchmarks and the proposed EquiContact are listed in Table I. We conducted a series of benchmark tests to demonstrate the effectiveness of each component in the proposed approach. The overall results are summarized in Table I.
TABLE I: Success rates of the insertion policies in real-world experiments for the proposed and benchmark approaches. “In-Dist.” denotes in-distribution data and “OOD” denotes out-of-distribution data. For the In-Dist. (in distribution) scenario, the initial pose of the end-effector is randomized around the flat platform.
Methods Observation Action Cameras Test Scenario Success Rate
ACT w/o GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 2 / 10
ACT w/ GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
CompACT [World Pose, FT] [World Pose, Gains] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
Flat Platform (OOD) 0 / 10
EquiContact [GCEV, FT] [Relative Pose, Gains] 2 Wrists Cameras Flat Platform (OOD) 20 / 20
(Place, Ours) Tilted Platform ($30^{\circ}$, OOD) 19 / 20
V-A1 Demonstration of Compliance
We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC). Results for this comparison are shown in the 1st and 2nd rows of Table I. Without GAC, the ACT model shows significantly lower success rates. The failure mode of the ACT w/o GAC involves collision: as the robot approaches the platform, excessive contact forces trigger safety shutdowns, preventing the task from completing. Due to the safety issue, we limited the number of trials without GAC to 10. This result demonstrates that the presence of compliant control is almost a decision factor between success and failure for contact-rich tasks.
To further analyze the effects of compliant gain control, we compare ACT with GAC and CompACT under in-distribution (In-Dist.) platform settings. Although both methods achieve perfect success rates, their force profiles during insertion differ substantially. As shown in Fig. 4, CompACT, which outputs task-adaptive admittance gains based on force-torque feedback, consistently produces lower interaction forces in all directions. The effectiveness of the CompACT compared to the baseline ACT was already presented in [17].
Figure 4: Force profiles of CompACT and ACT with GAC (fixed gains) during insertion tasks are presented. The CompACT with force-torque sensor inputs and output gains shows lower interaction force in all directions.
V-A2 Demonstration of Equivariance
Although the CompACT succeeds in insertion tasks in trained scenarios without excessive force exertion, it fails to generalize to spatially unseen configurations. This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance. The result of applying CompACT to the translationally unseen cases is shown in the Table I $3$rd row - Flat Platform (OOD). Note that the flat platform is only utilized, meaning it is only randomized translationally. We tested for $10$ cases and did not conduct more tests because it resulted in a $0\%$ success rate.
In contrast, the proposed method (EquiContact) achieves perfect success rates on the translationally unseen flat platform, as can be seen in the $4$th row of Table I. As shown in Section IV, EquiContact has $SE(3)$ vision-to-force equivariance, achieving a near-perfect success rate, even on the tilted platform, which undergoes a full $SE(3)$ transformation. We attribute the single failure case to a large error from Diff-EDFs that exceeded the noise level applied during training.
TABLE II: Success Rates of the proposed EquiContact for a full pipeline of pick and place.
Test Scenario Success Rate Failure Cases
Flat Platform (OOD) 20 / 20 N/A
Tilted Platform ($30^{\circ}$, OOD) 18 / 20 1 Pick, 1 Place
The result of EquiContact for the full pick-and-place task is summarized in Table II. The EquiContact also demonstrates a near-perfect success rate in the full pick-and-place pipeline for peg-in-hole tasks. However, as the whole task is formulated in sequential stages, the error in the previous stage tends to propagate to the following stages, leading to slightly increased failure cases.
From these experimental results, we draw the following key takeaway.
Remark 1 (Anchoring a Localized Policy on the Reference Frame for Equivariance).
The G-CompACT policy $\pi_{\theta}$ operates exclusively on localized inputs: GCEV, FT values in the end-effector frame, and images from wrist cameras. Its outputs, relative poses, and admittance gains are likewise defined in the local end-effector frame. The localized policy is anchored to a reference frame generated by the Diff-EDF planner. This architectural design induces spatial generalization and $SE(3)$-equivariance. Under this principle, one may hypothesize the reason for the spatial generalizability of recent vision-language-action models, which utilize both wrist and external cameras. The wrist camera provides a localized policy, anchored in the reference frame obtained from the external cameras and proprioceptive information. In fact, learning a reference frame from proprioceptive information (e.g., joint position) and the external camera is subtle because of its black-box nature. In contrast, our method, which utilizes a point cloud-based reference frame approach, is more interpretable. Thus, we propose “anchoring localized policies on globally estimated reference frames” as a general framework for the divide-and-conquer philosophy to enhance spatial generalization and interpretability. In our case, this structure also enables provable $SE(3)$ equivariance via differential geometric design.
V-B Failure Cases and Performance Recovery via Data Augmentation
TABLE III: Success Rates of the G-CompACT (insertion) trained with the base dataset and augmented dataset for the flat platform under the presence of visual distractor scenario and tilted platform with large angle cases. The evaluation is conducted with the ground-truth reference frames.
Test Scenario Base Dataset Aug. Dataset
W/ Visual Distractor 4 / 10 9 / 10
Tilted Platform ($45^{\circ}$) 5 / 10 9 / 10
Although the EquiContact framework is vision-to-force equivariant in theory, this guarantee holds only under Assumptions 1 and 2. In particular, Assumption 2, which requires left-invariance of the visual features, is more challenging to enforce in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias. As a result, the proposed G-CompACT algorithm shows degraded performance in scenarios with visual distractors or on a more severely tilted platform $(45^{\circ})$ – see Table III (“Base Dataset” column). Note that the larger tilting angle results in more out-of-distribution images, as the wrist cameras begin to capture more unseen background scenes, as illustrated in Fig. 3. To accommodate this, we augment the base dataset with $20$ demos collected using the visual distractor on a fixed, flat platform and $20$ demos collected from a $30^{\circ}$ angle of the tilted platform. After training with the augmented dataset, the success rates return to normal levels, as shown in the “Aug. Dataset” column of Table III, demonstrating the performance recovery achieved through data augmentation.
VI Conclusion
In this work, we introduced EquiContact, a vision-to-force equivariant policy for spatially generalizable contact-rich tasks. By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations. We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations and validated its effectiveness through real-world experiments on PiH benchmarks. Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates. Through extensive benchmark studies, we highlighted the effectiveness of the three principles – compliance, localized policy, and induced equivariance – for achieving spatial generalizability in contact-rich manipulation. We conclude that these principles offer a simple yet powerful design guideline for developing spatially generalizable and interpretable robotic policies complementing recent trends in end-to-end visuomotor learning and enabling a structured divide-and-conquer approach.
Supplementary Material
SI Implementation Details
In this section, we provide the details of the hardware code implementation. Since every hardware system has a different implementation, our goal is to provide a comprehensive overview and a guide to our code distribution. We will use the exact names of the nodes to enable readers to easily understand our codebase.
SI-A Implementation Pipeline
The pipeline visualized in Fig. S1 is an interconnected system of camera, robot, and controller nodes with two policy services, all implemented in ROS2 Humble. The two policies are implemented as services because they are activated infrequently and latencies in activating the policies initially do not affect the behavior in practice. The execution of the full pipeline is divided into two primary stages: querying Diffusion-EDF for initial pick and insertion end-effector poses, and then executing the appropriate ACT sub-policies for these two steps after reaching a neighborhood of the predicted poses. In the first stage, a request to the Diffusion-EDF Service is made through the FullPipelineClient, which prompts the service to pull the individual point clouds published by the Orbbec Node. The service subsequently merges the point cloud in the spatial frame, preprocesses it, and passes it to the initialized Diff-EDF model to produce both pick and insertion poses. The Diff-EDF model, as covered in the following section, is formulated as a “double-pick” model that assumes that a fixed grip will be executed, and a cached empty gripper point cloud is passed along with the scene point cloud to get both predictions.
In the second stage, the FullPipelineClient requests the G-CompACT Service node to complete the task with these two poses as noisy references. The service initializes both pick and place policies and executes the tasks serially, raising flags to signal completion of each step. For raw camera images and proprioceptive information, the service subscribes to nodes for each camera (from Wrist Camera Nodes) and the Robot Node respectively; note that the controller and proprioceptive helper are encapsulated in one Robot Node that manages both publishing proprioceptive information and executing policy predictions. During the pick portion, the pick ACT model only publishes end-effector poses, and fixed high admittance gains are utilized because the demonstrations encourage the model to make limited contact with the peg. The pick model has a fixed time to complete the task before a grasp is attempted, and the robot moves to the next step. Similarly, the place ACT policy is executed given a fixed-time limit, but for the place model, the force-torque data is also passed as an input, and admittance gains are outputted along with the end-effector poses. A fixed-time limit for both phases is used because it is easily implemented and works more consistently than other, more complicated completion heuristics.
Figure S1: Whole pipeline implemented (via client named as FullPipelineClient) with ROS2 is presented.
SI-B Diffusion-EDF
SI-B1 Implementation Heuristics
Instead of following the original Diff-EDF pipeline, which utilizes pick-and-place models, we used two pick models. This decision mainly stems from the task setup, where the peg is upright, and the peg is grasped by the gripper in an aligned, upright pose. The core difference of the pick and place model is that the place model needs to get the grasp point cloud after each grasp to handle the right equivariance of the model. However, from task setup, we bypass this right equivariance issue, removing the necessity of the place model.
We also used the post-processing heuristics to filter the output pose of the Diff-EDF. The Diff-EDF outputs $20$ candidate target poses for picking and $10$ candidates for the place, which are ranked by the energy level. Although in theory, the lower energy poses should result in a better pose, we found out that this does not hold in practice. Instead, we figured out that the Diff-EDF have low variations on the position of the “tip”. To elaborate, suppose a picking policy returns the desired poses $g_{pick}^{i}=(p_{pick}^{i},R_{pick}^{i})$, which are the end-effector poses, not the tip, or Tool-Center-Point (TCP) poses. Then, the position of the tip estimated by Diff-EDF can be calculated via
$p_{tip,pick}^{i}=p_{pick}^{i}+R_{pick}^{i}p_{et,pick}$
where $p_{et,pick}\in\mathbb{R}^{3}$ denotes the coordinate of the gripper’s tip with respect to the end-effector body frame. Using this property and the task setup that the peg is always upright, which gives $R_{peg}$, we recalculate the Diff-EDF of the picking as
$p_{pick}=\bar{p}_{tip,pick}-R_{peg}^{T}p_{et,pick}$
where $\bar{p}_{tip,pick}$ is the mean of $p_{tip,pick}^{i}$ for $i=1,\cdots,N$. On the other hand, for placement, we did not make any assumptions about the orientation of the hole-platform structure. The heuristic we used for the place is as follows:
$p_{place}=\bar{p}_{tip,place}-\bar{R}_{place}^{T}p_{et,place},$
where $\bar{R}_{place}$ is the averaged rotation of $R_{place}^{i}$, and $\bar{p}_{tip,place}$ is the averaged position of the tip calculated similarly from the picking case.
SI-B2 Accuracy of the Diff-EDF
Here, we present detailed error values of the Diff-EDF evaluated on the training dataset. We note that we followed the implementation pipeline of the original Diff-EDF package, but just utilized the two pick models as described in Section SI-B1. The resulting root mean squared error (RMSE) values are shown in Table S1. The rotational errors are calculated via $R_{e}=R_{gt}R_{pick}^{T}$ and converted to $xyz$ euler angle representation, where $pick$ is replaced by $place$ for the placing task.
TABLE S1: RMSE error values of Diff-EDFs. The dimension of translational error is $\mathrm{mm}$ and the rotational error is $\mathrm{deg}$.
Pick Place
Translation $x$ ($\mathrm{mm}$) 16.76 4.981
Translation $y$ ($\mathrm{mm}$) 5.886 7.422
Translation $z$ ($\mathrm{mm}$) 10.79 5.236
Rotation $x$ ($\mathrm{deg}$) 19.88 13.74
Rotation $y$ ($\mathrm{deg}$) 40.55 18.99
Rotation $z$ ($\mathrm{deg}$) 106.7 91.41
As mentioned in Section I, the error in translation ranges from $5\mathrm{mm}$ to $15\mathrm{mm}$, and rotational error can range up to $40^{\circ}$. We note that the $z$-directional rotation error can be arbitrarily high in our case because our PiH task is symmetric.
SI-C Tips on applying Geometric Admittance Control
Here, we present a few tips on the application of geometric admittance control. The GAC is implemented as a part of a ROS2 node, running with $200\mathrm{Hz}$.
SI-C1 ROS2 Implementation
The Robot Node acts as the central controller, receiving and executing desired pose commands using a Geometric Admittance Control (GAC). Its main execution loop operates at $200\mathrm{Hz}$, ensuring smooth and compliant motion in response to external interactions. The node subscribes to the “action” topic, which contains the pose command and admittance gains command. Additionally, the controller subscribes to the “rebias” topic in order to preserve the accuracy of force readings. When a message on this topic is received, a callback raises a flag, triggering a re-biasing procedure for the FT sensor during the next control cycle. Within its main loop, the controller sends the intended pose and processed FT sensor data to the GAC function, which generates the compliant commands for the robot’s low-level motion controller.
SI-C2 Handling Force-Torque Sensor Data
The first consideration is how to handle the Force-Torque (FT) sensor attached to the end-effector. It is always favorable to subtract the bias of the FT sensor after finishing each task, such as picking up the peg. We collect the FT sensor value for $2$ seconds and take a mean value to use as a bias. The other is to filter the FT sensor. Since the FT sensor is noisy, it is strongly recommended to implement a low-pass filter. However, having a low cut-off frequency can lead to serious phase lag, resulting in so-called “bouncing” behavior. We designed a simple 1st-order low-pass filter and implemented it using a state-space representation in discrete time.
SI-C3 Implementation Tips
As mentioned in Section. IV, the detailed implementation method depends on the low-level servo controller type. Typical controllers provided by robot manufacturers are joint position control, joint velocity control, end-effector position control, and end-effector velocity control. If a torque controller is readily available, then directly implementing a geometric impedance controller (GIC) is recommended, but joint friction compensation needs to be handled appropriately. The end-effector velocity control can be straightforwardly applied from (IV-B1), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command. One caveat here is the $V^{b}$, which is body-frame end-effector velocity. The common velocity provided by the manufacturer is defined on the Cartesian (Spatial) frame. If end-effector velocity is only offered in the Cartesian frame $V\in\mathbb{R}^{6}$, then
$V^{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}V,$
where $R$ is a rotation matrix from the current configuration. However, sometimes, it is provided in the Cartesian frame for translational velocity and the body frame for rotational velocity. The users should verify the frames of the velocities by running the robot in one direction and logging its values. For the joint space-based approach, one can use
$\begin{split}\dot{q}_{d}(k)\!&=\!\dot{q}(k)\!+\!T_{s}\cdot J_{b}(k)^{-1}M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),\\ q_{d}(k)\!&=\!q(k)+T_{s}\cdot\dot{q}_{d}(k),\end{split}$
where $J_{b}$ denotes a body-frame Jacobian matrix. $\dot{q}_{d}(k)$ is provided to joint-velocity controller and $q_{d}(k)$ is provided to the joint-position controller. The Jacobian matrix also needs to be identified concerning the end-effector velocity, i.e., if the end-effector velocity is given in the Cartesian frame, then the Jacobian matrix $J$ is also provided in the Cartesian frame. In this case, the body-frame Jacobian matrix can be obtained via
$J_{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}J$
When the Jacobian matrix is not provided, an alternative is to use the robotics toolbox from Python or Matlab. First, use the DH parameters provided by the manufacturer (or URDF), then build the Jacobian matrix function. The function can be cast into a lambda function in Python. For Python, the Jacobian matrix function can also be wrapped into executable .so file using Matlab code builder and executed in the loop for fast computation (well below $1\mathrm{ms}$).
SI-D Teleoperation Details
The architecture of our teleoperation system is explained to make it easier to reproduce and adapt our work. The system is constructed modularly, with the human-operator interface decoupled from the robot controller. This allows greater adaptation to different robotic platforms and control algorithms. Two ROS nodes make up the majority of the architecture: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
The teleoperation node acts as the principal interface between the human operator and the robot controller. A 6-DOF SpaceMouse is used to record input, which is then converted into the necessary end-effector pose. The main teleop function of this node operates at a frequency of $200\mathrm{Hz}$ inside a timer loop. It adjusts the desired end-effector pose incrementally during each cycle by reading velocity commands from the SpaceMouse.
The SpaceMouse provides velocity commands $V_{sm}=[v_{sm}^{T},\omega_{sm}^{T}]^{T}\in\mathbb{R}^{6}$, where $v_{sm}\in\mathbb{R}^{3}$ is translational velocity and $\omega_{sm}\in\mathbb{R}^{3}$ is rotational velocity. The change in desired position, $\Delta p$, is computed by scaling $v_{sm}$ with a position offset:
$\Delta p=v_{sm}\cdot\mathtt{pos\_offset}$
The target position is then updated as:
$p_{d}(k+1)=p_{d}(k)+\Delta p$
The desired end-effector orientation, $R_{d}(k+1)$, is updated by composing the previous rotation with the incremental rotation matrix, $\Delta R$. The small, incremental angles, $\Delta\theta$, are computed by scaling $w_{sm}$ with rotation offset:
$\Delta\theta=w_{sm}\cdot\mathtt{rot\_offset}$
The $\Delta\theta$ is then converted to $\Delta R$ using Euler angle as $\Delta R=\mathtt{euler2rotmat(\Delta\theta)}$. The target orientation is then updated as:
$R_{d}(k+1)=R_{d}(k)\cdot\Delta R$
This calculated pose is then published as a command for the robot controller at a frequency of $200\mathrm{Hz}$. To create a dataset of the teleoperation commands, a separate timer logs the same message at $30\mathrm{Hz}$. This implements a teleoperation that operates on the robot base frame for translational movement and the end-effector frame for rotational movement.
This teleoperation node was created with multiple operation modes to improve usability. For end-effector alignment with tilted platforms, the operator can switch the SpaceMouse’s control frame between the end-effector frame and the robot’s base frame. Additionally, a dedicated button allows users to modify the sensitivity offsets for position and rotation, enabling users to alternate between a translation-focused and an orientation-focused mode. The operator can globally scale motion sensitivity by cycling through low, medium, and high speed levels. The system also includes specific features, such as an “insertion mode” that operates in the end-effector frame for the translational movement to help with high-precision alignment tasks, as well as requests to set the robot’s end effector at “home” or a randomized pose.
Furthermore, the operator can dynamically adjust the parameters of the Geometric Admittance Controller via dedicated keyboard inputs. There are four gain profiles: Free mode applies high stiffness gains $(1000)$ to all axes for precise motion in open space; Contact mode uses low z-axis gain $(300)$ with high gain elsewhere $(1500)$ to exert less force on surfaces; Insertion mode reverses this, with high z-axis gain $(1500)$ for precision while allowing compliance in other directions $(300)$; and Compliant mode sets all gains low $(300)$ to make the end-effector highly responsive to external forces. These gain profiles are published as action messages, together with poses and gripper states.
SI-E G-CompACT Training Details
During the training of the G-CompACT, we added the noise to the known fixed pose of the hole and peg. As mentioned earlier, this is because the output of the Diff-EDFs $g_{{}_{EDF}}$ is biased. By adding the noise to the goal pose signal, we encourage the model to focus on the image of the wrist camera instead of the incorrect GCEV $e_{{}_{G}}$ signal. For our training, we added noise to every step of the demonstration dataset drawn from a uniform distribution. The noisy target poses $\tilde{g}_{ref}=(\tilde{p}_{ref},\tilde{R}_{ref})$are calculated as
$\begin{split}\tilde{p}_{ref}&=p_{ref}+n_{p},\\ \tilde{R}_{ref}&=R_{ref}\cdot\mathtt{euler2rotmat}(n_{r}),\end{split}$
where $n_{p}\sim\mathcal{U}(-0.02,0.02)\in\mathbb{R}^{3}$ and $n_{r}\sim\mathcal{U}(-8^{\circ},8^{\circ})\in\mathbb{R}^{3}$, which denotes $2\mathrm{cm}$ of translational and $8^{\circ}$ (degree) of rotational error.
To encode the element on $SE(3)$ for learning, we used position & rotation vector representation for G-CompACT. However, in benchmark models of CompACT and ACTs, we utilized position & rot6d representation. This is because the rotation vector representation has a sign flip issue when the rotation angle is near $180^{\circ}$, but the rot6d representation does not have that problem. For G-CompACT, as the actions are represented with respect to the current end-effector frame, the transformation angles are relatively small and do not have a sign flip problem.

### v2

@@
 We focus on achieving robust spatial generalization of the policy for the peg-in-hole (PiH) task trained from a small number of demonstrations.
-We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant ACT, G-CompACT).
+We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant Action Chunking Transformers, G-CompACT).
 G-CompACT operates using only localized observations (geometrically consistent error vectors (GCEV), force-torque readings, and wrist-mounted RGB images) and produces actions defined in the end-effector frame.
@@
 We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance.
-Real-world experiments on PiH tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles.
+Real-world experiments on PiH, screwing, and surface wiping tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles.
 The experimental videos can be found on the project website: https://sites.google.com/berkeley.edu/equicontact
@@
 This has led to massive efforts to build large datasets [1] for training policies with general knowledge.
-However, such policies often lack spatial generalizability and therefore require a large amount of data to learn robust policies.
-As described in [9], both action chunking transformers (ACT) [6] and diffusion policy (DP) [5] exhibit near-linear performance growth as the demonstration dataset size increases, suggesting that the trained policies do not inherently generalize well to new spatial configurations.
+However, such policies often lack spatial generalizability and therefore require a large amount of data to learn robust behaviors.
+As described in [9], both action chunking transformers (ACT) [6] and diffusion policy (DP) [5] are evaluated only within the limited spatial variations.
+Furthermore, both methods exhibit near-linear performance growth as the demonstration dataset size increases, suggesting that the trained policies do not inherently generalize well to new spatial configurations, but rather tend to interpolate between seen demonstrations.
 An alternative line of recent research focuses on leveraging symmetry—particularly equivariance—to enhance spatial generalizability, thereby improving sample efficiency during training [10, 11].
@@
 As a result, it becomes more attractive for users to use standard models trained with massive datasets in many instances.
-[…]
-In this paper, we propose EquiContact, a hierarchical SE(3) vision-to-force equivariant policy for spatially generalizable contact-rich tasks.
-Our proposed method consists of two main components.
-The first is a high-level planner consisting of a Diff-EDF model [14], which uses global point cloud data to provide a local reference frame for placing the peg relative to the hole.
-The second, low-level compliant visuomotor policy is a variant of ACT [6], which we refer to as Geometric Compliant ACT (G-CompACT).
-This policy handles detailed motion and contact interaction using force-torque feedback and RGB inputs from wrist-mounted cameras.
-A key design feature of G-CompACT is that it only relies on local information—specifically, the force-torque signal in the end-effector frame, a geometrically consistent error vector (GCEV) [11], and wrist camera inputs.
-The output of G-CompACT is the desired pose and admittance gains, which are then sent to the geometric admittance controller (GAC) module to execute compliant control.
-We emphasize that our contribution is on the structural framework rather than the specific choice of algorithms.
-For example, one could replace the Diff-EDF with other equivariant methods, such as ET-SEED [15], or replace the ACT with the DP or their variants, e.g., Diffusion Transformer (DiT) [16].
+In [11], a $SE(3)$-equivariant gain-scheduling policy using geometric impedance control (GIC) [12, 13] was proposed to solve peg-in-hole (PiH) problems.
+Inspired by the view that many manipulation tasks can be framed as pick-and-place problems [4], we modeled PiH as a compliant pick-and-place task, where final peg poses are provided by vision-based $SE(3)$-equivariant models such as Diffusion-EDF (Diff-EDF) [14].
+Since both the high-level planner and low-level variable impedance controller are equivariant, they can be combined to form a vision-to-force equivariant policy.
+However, in practice, Diff-EDF’s placement accuracy proved insufficient for tight insertion tasks, which require sub-millimeter precision (details in Appendix -A).
+This revealed a key limitation: high-level vision planners may capture global structures but struggle with precision and contact-sensitive execution.
+From this insight, we introduce an intermediate layer between the planner and the low-level controller, which provides real-time visual feedback to correct the residual errors of the high-level planner.
+In this paper, we propose EquiContact, a hierarchical $SE(3)$ vision-to-force equivariant policy for spatially generalizable, contact-rich tasks.
+It consists of two main components: a high-level planner using Diffusion Equivariant Descriptor Fields (Diff-EDF) [14], which estimates a local reference frame from point clouds, and a low-level compliant visuomotor policy based on Action Chunking Transformer (ACT) [6], which we refer to as Geometric Compliant ACT (G-CompACT).
+A key design feature of G-CompACT is that it only relies on local information: the force-torque signal in the end-effector frame, a geometrically consistent error vector (GCEV) [11], and wrist camera inputs.
+The output of G-CompACT is the relative desired pose and admittance gains, which are then sent to the geometric admittance controller (GAC) module to execute compliant control.
+Our contribution lies in the framework design, not in specific model choices; for example, Diff-EDF could be replaced by ET-SEED [15], or ACT by other visuomotor policies.
 The main contributions of this paper are as follows:
@@
   2. 2.
-    We demonstrate that EquiContact achieves near-perfect success rates and spatial generalizability in real robot experiments involving tight insertions.
+    We identify three key principles for spatially generalizable contact-rich manipulation: left-invariant compliant control action (via GAC [11]), localized policy (left invariance), and induced equivariance.
+These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [10].
   3. 3.
-    We identify three key principles for spatially generalizable contact-rich manipulation: Compliance, Localized Policy (Left invariance), and Induced Equivariance.
-These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [10].
+    Under these principles, we present the necessary conditions for $SE(3)$ vision-to-force equivariant policy, and mathematically prove the equivariance property of EquiContact.
+  4. 4.
+    We demonstrate that EquiContact achieves near-perfect success rates and spatial generalizability when these conditions are met in real robot experiments involving peg-in-hole, screwing, and surface wiping tasks.
 From these key principles, we propose a general framework to enhance the spatial generalization and interpretability of vision-based policies, namely, “anchoring localized policy on globally estimated reference frame.” We emphasize that our work provides complementary insights to recent trends of robot learning [1, 8, 16, 2] that aim to build generalist policies from a large-scale demo dataset.
@@
 We have significantly improved CompACT by incorporating a provable $SE(3)$ equivariant structure.
-Equivariant Methods Equivariant methods aim to leverage the inherent symmetries of robot tasks, typically modeled as group transformations such as $SE(3)$, to improve the sample efficiency of imitation learning.
-Early approaches, such as Neural Descriptor Fields (NDFs) [22] and Transporter Networks [3], incorporated equivariance to $SE(3)$ and $SO(2)$ transformations, utilizing scene point clouds and top-down views, respectively.
-Equivariant Descriptor Fields (EDFs) [23] reformulated the NDF approach into a fully end-to-end learning method with $SE(3)$ bi-equivariance property, which is further improved by Diffusion-EDF (Diff-EDF) [14] to improve training and inference performance.
-Other equivariant models that extract desired keyframes from point clouds include [24].
-3D Equivariant extensions of DP [25, 26, 15] and Flow-Matching [27, 28] policies have been explored in recent literature and combine $SE(3)$ equivariant observation encoders with appropriately modified diffusion/flow-matching processes.
-Compared to recent works of [29, 30], we generalize the table-top manipulation to contact-rich tasks while achieving full $SE(3)$ equivariance from vision to force/control level, bridging vision, control, and force interaction under a unified framework.
+Equivariant Methods Earlier equivariant approaches attempted to handle manipulation tasks as an extension of pick-and-place tasks, by leveraging $SE(3)$ equivariance from point clouds [22, 15, 14] or $SO(2)$ equivariance [3] from top-down views.
+Equivariant approaches have been extended to visuomotor policies, such as DP or flow matching, [23, 24], using point clouds.
+[25, 26] proposed $SO(2)$ equivariant visuomotor policies using 2D images, not fully considering $SE(3)$.
+In contrast, our approach induces full $SE(3)$ equivariance from vision to control force without relying on explicitly equivariant neural networks, but using structured observations and actions via geometrical canonicalization.
+Furthermore, by integrating with $SE(3)$ equivariant control, we generalize beyond table-top settings to contact-rich manipulation.
+Manipulation in Object Frame Our induced $SE(3)$ equivariant approach relies on representing the visuomotor policy in the end-effector frame.
+While recent works [27, 28, 29] define policies in the target (object) frame, policy representation in the end-effector frame offers improved fidelity and robustness.
+This is because the estimated object frame can be noisy, and the end-effector frame is reliably obtained via forward kinematics.
+Importantly, compared to [27, 28], we explicitly link the choice of reference frame to the equivariance property, and unlike [29], which only handles translational transformations, our method can deal with full $SE(3)$ transformations of the reference frame.
 Figure 1: We propose an EquiContact, a hierarchical, provably $SE(3)$ vision-to-force equivariant policy for spatially generalizable contact-rich tasks.
-The proposed EquiContact consists of G-CompACT and Diffusion-EDF.
-The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen scenarios during evaluation.
+The proposed EquiContact consists of G-CompACT (Section.
+IV-B) and Diffusion-EDF (Section.
+IV-A).
+The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen task transformations during evaluation.
+The overall EquiContact algorithm is summarized in Algorithm 1.
 III Problem Definition
-In this paper, we aim to identify the key structural components of a policy capable of executing spatially generalizable, contact-rich manipulation tasks.
-As mentioned earlier, our proposed building blocks to achieve $SE(3)$ vision-to-force equivariant policy are: 1.
-Compliant Control, 2.
-Localized Policy, and 3.
-Induced Equivariance - which we will experimentally validate through specific data collection and evaluation setups, as detailed below.
-III-A Problem Setup
-Our primary focus is to provide a learning-based solution for a peg-in-hole task,a classic representative of contact-rich, force-based robotic manipulation and assembly tasks, using vision.
-Unlike prior work such as [11], which assumes that the desired hole pose is known and that the robot starts with a pre-grasped peg, we consider a more general scenario.
-The robot first needs to pick up the peg and then insert it into the hole, as shown in Fig.
-2, relying solely on vision and proprioception.
+In this paper, we aim to identify the key structural components required for learning policies that generalize spatially in contact-rich manipulation tasks.
+We will first focus on the peg-in-hole (PiH) problem as a representative force-based assembly task and validate the feasibility of the proposed approach to other contact-rich tasks later.
+Our proposed framework achieves $SE(3)$ vision-to-force equivariance through three essential design principles: (1) left-invariant compliant control, (2) localized policy, and (3) induced equivariance.
+These principles are validated through specific data collection and evaluation setups, as detailed below.
+Unlike prior work [11] that assumes a known hole pose and a pre-grasped peg, we consider a more general setup: the robot must first grasp the peg and then perform insertion using only vision and proprioception, as illustrated in Fig.
+2.
 We assume only that the peg is upright in the scene.
-Following the reasoning that every manipulation task can be regarded as a sequence of pick-and-place, we formulated peg-in-hole insertion as a compliant pick-and-place task.
-However, as highlighted in the earlier discussion, naively using the pick-and-place approach is insufficient because the high-level vision planner often lacks the accuracy required for precise and contact-rich tasks, which is $1\mathrm{mm}$ in our case (see Fig.
-2).
-To accommodate this, we propose a low-level policy that:
-  • provides real-time visual feedback to correct errors from the coarse high-level planner,
+Inspired by the idea that many manipulation tasks can be viewed as sequences of pick-and-place actions, we formulate insertion as a compliant pick-and-place task.
+However, high-level vision planners often lack the precision needed for tight-tolerance tasks like PiH, which in our case requires $1\mathrm{mm}$ accuracy.
+To accommodate this, we propose a low-level compliant policy that:
+  • provides real-time visual feedback to refine the coarse high-level command,
   • handles fine force-based interactions through compliance,
   • and achieves provable $SE(3)$-equivariance.
-Assuming that a high-level vision planner that returns approximate desired end-effector pick-and-place poses is readily available, we focus on developing equivariant compliant placing, i.e., an insertion policy via imitation learning, and regard picking as a spatial case of placing.
-To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location.
-This enables the policy to perform nearly perfectly in the trained scenario.
+We assume that a high-level vision planner (e.g., Diff-EDF) can generate approximate pick-and-place poses.
+Our focus is on developing an equivariant compliant placing policy, i.e., insertion policy, using imitation learning.
+We further assume the peg is approximately aligned with the gripper during placement, since arbitrary peg poses introduce two challenges: (1) imprecise grasps can lead to slippage during contact, and (2) compensating for slippage requires continuous estimation of the gripper-to-peg transformation, which is difficult to achieve reliably in real time.
+To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location, where its objective is to train a policy that performs nearly perfectly in the trained scenario.
 We then evaluate benchmark and proposed methods—trained solely on these limited demonstrations—across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
 Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented.
-$2$ external cameras with calibrated extrinsics and $2$ wrists cameras are installed.
+22 external cameras with calibrated extrinsics and 22 wrists cameras are installed.
 (Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance.
@@
 IV Solution Approach
-We now describe our proposed EquiContact policy, which integrates a high-level vision-based planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant control Action Chunking Transformer (G-CompACT).
-G-CompACT itself consists of a behavior cloning module built on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC).
-As discussed in Section III, because of space constraints, we focus on the insertion task in this paper, which will later be extended to include the picking task to complete the full pipeline implementation.
-We note that the Diff-EDF and G-CompACT are trained separately but are combined as a whole pipeline during inference time, as shown in Fig.
+We now describe our proposed EquiContact policy, which integrates a high-level vision planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant ACT (G-CompACT).
+G-CompACT consists of a behavior cloning module based on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC).
+As discussed in Section III, we focus on the insertion task, with extension to picking addressed later in the paper.
+The Diff-EDF and G-CompACT are trained separately and combined at inference, as shown in Fig.
 1.
+Conceptually, EquiContact follows a simple yet powerful principle: “anchoring a localized policy on a globally estimated reference frame.” In our framework, G-CompACT serves as the localized low-level policy, operating solely on observations defined in the current end-effector frame; wrist camera RGB images, force-torque (FT) signals, and a geometrically consistent error vector (GCEV) $e_{{}_{G}}$.
+It outputs the desired relative pose and admittance gains, also in the end-effector frame.
+As both inputs and outputs are defined locally, G-CompACT functions as a fully localized policy.
+The high-level planner, Diff-EDF, estimates the pose of the target (e.g., hole) in the world frame.
+At the inference, the robot moves near the estimated reference frame, and G-CompACT is activated to perform compliant motion using only local feedback.
+Because the low-level policy does not depend on absolute global inputs, it can transfer robustly to unseen spatial configurations when the estimated reference frame is provided.
+This divide-and-conquer design provides a general framework for enhancing both spatial generalization and policy interpretability for contact-rich, and more broadly, general manipulation tasks.
+In the remainder of this section, we formalize this spatial generalization property as $SE(3)$ equivariance and show how EquiContact satisfies this by construction through a rigorous mathematical derivation.
+We note that $\theta$, $\theta_{i}$, and $\phi$ with $i=1,2,3$ denote a learnable parameter of the neural network, without loss of generality.
+We also note that we follow the notation and definition conventions of [11, 30].
 IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
@@
 $g_{{}_{EDF}}=f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$ (1)
-where the neural network $f_{\theta_{1}}$ implements the mapping. Due to the left-equivariant property, it follows that:
-$g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
+where the neural network $f_{\theta_{1}}$ implements the mapping.
+Due to the left-equivariant property, it follows that: $g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
 Let $\mathcal{O}^{ref}\subset\mathcal{O}^{scene}$ denote the subset of points corresponding to the object of interest (e.g., the hole structure, see also Fig.
@@
 This encourages the model to focus on $\mathcal{O}^{ref}$, not the rest of the objects in the scene.
-While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and have high robustness to left $SE(3)$ transformations.
+While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and has high robustness to left $SE(3)$ transformations.
 To train Diff-EDF, the scene and grasp point clouds are collected together with the target reference frames, which represent the desired poses of the end-effector for pick-and-place operations.
@@
 The training process of Diff-EDF follows the procedure in [14].
-While the original Diff-EDF supports right-equivariance to handle transformations between the gripper and the grasped object, we did not utilize this feature.
-In practice, we observed that arbitrary peg transformations relative to the gripper introduced additional challenges.
-First, grasping the peg based on an arbitrary transformation leads to imprecise grasps, resulting in object slippage during contact.
-Second, the gripper-object rotation must be estimated continuously in real-time for G-CompACT to work under slippage.
-Instead, we enforced a consistent grasp orientation and relied only on the left-equivariance property.
-We refer the readers to [14, 10] for full details of Diff-EDF and to Supplementary Material  SI for implementation and heuristics.
-We emphasize again that Diff-EDF can be replaced by any point cloud-based method that outputs an approximate goal pose for (compliant) placing.
+Note that arbitrary peg transformations relative to the gripper can be handled via the right-equivariance feature of the Diff-EDF; however, due to the complications mentioned in Section.
+III, we enforced a consistent grasp and only utilized the left-invariance property.
+We refer the readers to [14, 10] for full details of Diff-EDF.
 IV-B Geometric Compliant control Action Chunking with Transformers (G-CompACT)
-IV-B1 Geometric Admittance Control (GAC)
-We implement the geometric impedance control (GIC) proposed in [12, 13] in the geometric admittance control (GAC) setup [11].
-We first note that we follow the notation used in [11].
-Let the end-effector pose be denoted as $g\in SE(3)$ in a homogeneous matrix representation, or simply $g=(p,R)$, where $p\in\mathbb{R}^{3}$ is a position of the end-effector and $R\in SO(3)$ is a rotation matrix of the end-effector.
-Given the desired end-effector pose $g_{d}=(p_{d},R_{d})$, the desired end-effector dynamics for the GAC setup is given as follows:
-$M\dot{V}^{b}+K_{d}V^{b}+f_{{}_{G}}=F_{e},$ (3)
-where $M\in\mathbb{R}^{6\times 6}$ is symmetric positive definite desired inertia matrix, $K_{d}\in\mathbb{R}^{6\times 6}$ symmetric positive definite damping matrix, $F_{e}\in\mathbb{R}^{6}$ is external wrench applied to the end-effector in end-effector body frame and $V^{b}\in\mathbb{R}^{6}$ is a body-frame end-effector velocity.
-Further, $f_{{}_{G}}=f_{{}_{G}}(g,g_{d},K_{p},K_{R})\in\mathbb{R}^{6}$ is a elastic wrench given by:
-$\begin{split}f_{{}_{G}}=\begin{bmatrix}f_{p}\\ f_{R}\end{bmatrix}=\begin{bmatrix}R^{T}R_{d}K_{p}R_{d}^{T}(p-p_{d})\\ (K_{R}R_{d}^{T}R-R^{T}R_{d}K_{R})^{\vee}\end{bmatrix},\end{split}$ (4)
-where $\hat{f}_{{}_{G}}\in se^{*}(3)$, with $K_{p},K_{R}\in\mathbb{R}^{3\times 3}$ symmetric positive stiffness matrices for the translational and rotational dynamics, respectively, and $(\cdot)^{\vee}$ denotes the vee-map.
-By a slight abuse of notation we will frequently denote as $f_{{}_{G}}\in se^{*}(3)$ to emphasize that it is an elastic wrench on $SE(3)$, since $f_{{}_{G}}$ is isomorphic to $\hat{f}_{{}_{G}}$ [13].
-Given the desired dynamics (3), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
-$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
-$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (5)
-where $T_{s}$ is a sampling time and $\hat{(\cdot)}$ denotes a hat-map.
-For the details of the GIC/GAC, we refer to [12, 11].
-The admittance control loop is implemented at a $200\mathrm{Hz}$ rate using ROS2.
-We refer to Supplementary Material SI for detailed implementation tips.
-IV-B2 G-CompACT
-Action Chunking with Transformer (ACT) is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks, originally introduced in [6].
-While the original ACT formulation encodes robot proprioception using joint positions, this representation lacks spatial equivariance, limiting generalization.
-Instead, we adopt a proprioceptive encoding based on the end-effector pose, following the approach in [17].
-In our previous work [11], we proposed a recipe for spatially equivariant policies, which involves a left invariance property and policy representation in the end-effector body frame.
-Following this principle and recent developments on relative action representations (similar to [29]), we structure our G-CompACT policy to take observations composed of: (1) Geometrically Consistent Error Vector (GCEV) proposed in [11], (2) FT sensor in the end-effector frame to capture contact behaviors, and (3) RGB images from wrist cameras.
-The model returns as action sequences: (1) relative pose from the current end-effector frame, and (2) admittance gains.
+IV-B1 G-CompACT
+G-CompACT is based on the Action Chunking with Transformer (ACT), which is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks [6].
+To make G-CompACT spatially equivariant, we follow the principles proposed in [11]: left-invariant policy and policy representation in the end-effector body frame.
+To achieve this, we structure the observations and actions of the G-CompACT as described below.
+First, the observations are given by, (1) Geometrically Consistent Error Vector (GCEV) $e_{{}_{G}}$ proposed in [11], (2) FT sensor in the end-effector frame $F_{e}$ to capture contact behaviors, and (3) RGB images $I_{w}=\{I_{w,1},I_{w,2}\}$ from wrist cameras (see Fig.
+3).
+For actions, we choose (1) relative pose from the current end-effector frame $g_{rel}$, and (2) admittance gains for Geometric Admittance Control (GAC) $(K_{p},K_{R})$.
+The details of GAC and the definition of gains will be provided later in the Section.
+Formally, let the G-CompACT method be $\pi_{\theta_{2}}$ such that
+$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq(g_{rel},K_{p},K_{R})(k),\quad o(k)\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (3)
+where $a(k)$ denotes the actions, and $o(k)$ denotes the observation at time step $k$.
+In fact, our G-CompACT also outputs the actions of chunk size $N$ following the standard action-chunking strategy.
+However, for the compact presentation, we will only consider the single-step action after undergoing proper processing, such as a temporal ensemble.
 The GCEV $e_{{}_{G}}(g,g_{{}_{EDF}})$ is defined as
 $e_{{}_{G}}(g,g_{{}_{EDF}})=\begin{bmatrix}R^{T}(p-p_{{}_{EDF}})\\ (R_{{}_{EDF}}^{T}R-R^{T}R_{{}_{EDF}})^{\vee}\end{bmatrix}$, (4)
-where $g=(p,R)\in SE(3)$ is a current end-effector pose and $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF.
-The derivation of the GCEV $e_{{}_{G}}$ is presented in [12].
-We note that the GCEV denotes an error vector on the end-effector body frame as emphasized in [11].
-The model outputs a chunk of $N$ relative pose actions:
-$\bar{g}_{d}(k+i)=g(k)\cdot g_{rel}(k+i),\quad i=1,\cdots,N,$ (7)
-where each $g_{rel}(k+i)\in SE(3)$ is the $i$th relative pose from the ACT model and $\bar{g}_{d}(k+i)$ is the $i$th temporary desired pose signal.
-A temporal ensembling mechanism is then applied to $\bar{g}_{d}(k+i)$ to compute the desired pose $g_{d}(k)$ for GAC.
-To respect the manifold structure of $SE(3)$, we use the scipy implementation for weighted averaging on the rotation matrix to get $R_{d}(k)$ and standard weighted averaging on positions for $p_{d}(k)$ separately.
-Our proposed G-CompACT also outputs the admittance gains matrices $K_{p}$ and $K_{R}$ in the GAC elastic wrench in Eq.
-(4).
-Unlike CompACT, the gain matrices are represented in the desired frame of the end-effector [13], removing the necessity of utilizing full-gain matrix representations and coordinate transformations to deal with spatial equivariance.
-Therefore, a diagonal matrix representation in $K_{p}$ and $K_{R}$ is sufficient to conduct the PiH task successfully, without requiring a Cholesky factorization to represent the gains, as in CompACT.
-Since the diagonal component representation of matrices is equivalent to a vector, the standard temporal ensemble can be applied.
-Similar to the relative pose action, the ACT provides temporary gain signals $\bar{K}_{p}$ and $\bar{K}_{R}$, followed by a temporal ensemble to obtain $K_{p}$ and $K_{R}$.
-Formally, let the G-CompACT method be $\pi_{\theta}$ such that
-$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq\{(g_{rel}(k+i),\bar{K}_{p}(k+i),\bar{K}_{R}(k+i))\}_{i=1}^{N},\\ o(k)&\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (8)
-where $\bar{K}_{p}$ and $\bar{K}_{R}$ denote the temporary admittance gains, $a(k)$ denotes the actions and $o(k)$ denotes the observation at time step $k$, $e_{{}_{G}}$ is GCEV defined in (6), $F_{e}$ is force-torque sensor reading on the end-effector frame, and $I_{w}=\{I_{w,1},I_{w,2}\}$ where $I_{w,1}$ and $I_{w,2}$ denote the images from the wrist cameras #1 and #2, respectively (Fig.
-2).
+where $g=(p,R)\in SE(3)$ is a current end-effector pose, $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF, and $(\cdot)^{\vee}$ denotes the vee-map, a mapping from $so(3)$ (Lie algebra of $SO(3)$) to $\mathbb{R}^{3}$.
+The physical meaning of GCEV is an error vector in the current end-effector frame between the current end-effector pose and the reference frame, considering the $SE(3)$ manifold structure, i.e., non-Euclidean.
+For the details of GCEV $e_{{}_{G}}$, we refer to [11, 12].
 Since the images $I_{w}$ are fed to the transformer decoder after being passed to the visual encoder structure, one can further consider the G-CompACT as
@@
 Figure 3: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$.
-As the left group action is applied on the end-effector, the wrist cameras start to see not only the optical table but also the outer backgrounds.
-$\{s\}$ denotes a spatial frame.
+As the left group action is applied on the end-effector, the wrist cameras start to see backgrounds other than the optical table.
+$\{s\}$ denotes a spatial frame, i.e., robot base frame or world frame.
 The following proposition shows the left-invariance of the G-CompACT method in the end-effector frame.
@@
 $a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi_{\theta_{2}}(o(k)).$ (7)
+The proof is presented in Appendix. -B. In what follows, we present an equivariant property of the pose signal produced by G-CompACT when described in the spatial frame.
+Corollary 1 ($SE(3)$ Equivariance of G-CompACT).
+The G-CompACT $\pi_{\theta_{2}}$ represented in the spatial frame satisfies the following equivariance property:
+$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\pi_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (8)
 Proof.
-The left-translated observation signals $o(k)$ reads that:
-$g_{l}\circ o(k)=(g_{l}\circ e_{{}_{G}},g_{l}\circ F_{e},g_{l}\circ I_{w}).$ (12)
-As was shown in Lemma 1 of [11], the GCEV $e_{{}_{G}}$ is left invariant as
-$g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}})=e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}})=e_{{}_{G}}(g,g_{{}_{EDF}}).$
-The force-torque sensor values are left-invariant because they are already defined with respect to the end-effector frame [11], and the visual representation vectors satisfy left invariance due to Assumption 1.
-Combining all these properties, it follows that
-$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)=\pi_{\theta_{2}}(o(k)),$ (13)
-which shows the left invariance of the G-CompACT policy on the end-effector frame. ∎
-In what follows, we demonstrate that the pose signal produced by G-CompACT is equivariant when described in the spatial frame.
-Corollary 1 ($SE(3)$ (Left) Equivariance of G-CompACT).
-Let the output of G-CompACT in the spatial frame be $\bar{g}_{d}$ via (without considering the outputs related to gains)
-$\bar{g}_{d}=g\!\cdot\!\pi_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (14)
-Then,
-$g_{l}\bar{g}_{d}$ $=g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w})$
-$=g_{l}g\cdot\pi_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (15)
-Since the temporary desired pose on the spatial frame signal, $\bar{g}_{d}$, is $SE(3)$ equivariant, the temporal aggregated signal $g_{d}$ is also equivariant.
-Therefore, by combining the G-CompACT with a spatial frame representation and temporal ensemble, we will now define the final form of G-CompACT as $\hat{\pi}_{\theta_{2}}$, which is given by:
-$(g_{d},K_{p},K_{R})=\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w}).$ (16)
-and satisfies the following equivariance property:
-$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\hat{\pi}_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (17)
+Notice that the desired pose signal in the spatial frame $g_{d}$ is obtained via $g_{d}=gg_{rel}=g\pi_{\theta_{2}}(o(k))$.
+Then, utilizing the left-invariance property, $g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ o(k))=g_{l}g_{d}$.
+∎
 To train a G-CompACT, we collect the expert demonstration using teleoperation on the fixed pose of the platform.
@@
 The rest of the training follows the standard imitation learning pipeline.
+IV-B2 Geometric Admittance Control (GAC)
+We implement the geometric impedance control (GIC) proposed in [12, 13] in the geometric admittance control (GAC) setup [11].
+Let the end-effector pose be denoted as $g\in SE(3)$ in a homogeneous matrix representation, or simply $g=(p,R)$, where $p\in\mathbb{R}^{3}$ is a position of the end-effector and $R\in SO(3)$ is a rotation matrix of the end-effector.
+The GAC operates with the $(g_{d},K_{p},K_{R})$ signal calculated from G-CompACT, where the desired end-effector pose is calculated via $g_{d}=gg_{rel}$.
+Given $g_{d}=(p_{d},R_{d})$, the desired end-effector dynamics for the GAC setup is written as follows:
+$M\dot{V}^{b}+K_{d}V^{b}+f_{{}_{G}}=F_{e},$ (9)
+where $M\in\mathbb{R}^{6\times 6}$ is symmetric positive definite desired inertia matrix, $K_{d}\in\mathbb{R}^{6\times 6}$ symmetric positive definite damping matrix, $F_{e}\in\mathbb{R}^{6}$ is external wrench applied to the end-effector in end-effector body frame and $V^{b}\in\mathbb{R}^{6}$ is a body-frame end-effector velocity.
+$K_{d}$ matrix is selected to ensure overdamped system, as $K_{d}=3\cdot\text{blkdiag}(\sqrt{K_{p}},\sqrt{K_{R}})$.
+Further, $f_{{}_{G}}=f_{{}_{G}}(g,g_{d},K_{p},K_{R})\in\mathbb{R}^{6}$ is a elastic wrench given by:
+[…]
+where $K_{p},K_{R}\in\mathbb{R}^{3\times 3}$ symmetric being positive stiffness matrices for the translational and rotational dynamics, respectively.
+Following the standard admittance control formulation, the desired end-effector pose command is calculated using (9), which is passed to the lowest level tracking controller.
+For details on GIC/GAC, we refer readers to [12, 11].
+The admittance control loop is implemented at a $200\mathrm{Hz}$ rate using ROS2.
 IV-C EquiContact
@@
 The overall pipeline of the EquiContact is presented in Fig.
+1, and also summarized in Algorithm.
 1.
-Here, we demonstrate that the proposed EquiContact possesses the equivariance property.
-Let EquiContact be written as $h_{\theta}$ so that $h_{\theta}(g,g_{ref},F_{e})\mapsto f_{{}_{G}}$, i.e., $h_{\theta}:SE(3)\times SE(3)\times se^{*}(3)\to se^{*}(3)$.
+In Proposition 2, we demonstrate that the proposed EquiContact possesses the equivariance property.
+Let EquiContact be written as $h_{\theta}$ so that $h_{\theta}(g,g_{ref},F_{e})\mapsto f_{{}_{G}}$, i.e., $h_{\theta}:SE(3)\times SE(3)\times\mathbb{R}^{6}\to\mathbb{R}^{6}$.
 Proposition 2.
-The EquiContact Policy $h_{\theta}$ is equivariant if it is described relative to the spatial frame.
-Proof.
-Suppose that the Assumption 1 and 2 hold.
-Let the object of interest, e.g., a peg for the picking task and a hole for the placing task, be observed by $\mathcal{O}^{ref}$, $I_{w}$ with its pose given by $g_{ref}$, so that the left-translated $g_{l}\cdot g_{ref}$ is observed by $g_{l}\circ\mathcal{O}^{ref}$ from the point cloud, and $g_{l}\circ I_{w}$ by the left-translated end-effector attached wrist camera as described in Fig.
-3.
-First, notice that $h_{\theta}$ can be fully written as
-$h_{\theta}(g,g_{ref},F_{e})$ $=f_{{}_{G}}(g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,f_{\theta_{1}}(\mathcal{O}^{ref})),F_{e},I_{w}))$ (18)
-$=f_{{}_{G}}(g,\underset{=(g_{d},K_{p},K_{R})}{\underbrace{\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w})}}).$
-Then, when both $g$ and $g_{ref}$ undergoes a left transformation $g_{l}$, from Assumption 1 and Corollary 1, the following holds:
-$h_{\theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})$
-$=f_{{}_{G}}(g_{l}g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g_{l}g,f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref})),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
-$=f_{{}_{G}}(g_{l}g,\hat{\pi}_{\theta_{2}}(e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
-$=f_{{}_{G}}(g_{l}g,g_{l}g_{d},K_{p},K_{R})=f_{{}_{G}}(g,g_{d},K_{p},K_{R})$
-$=h_{\theta}(g,g_{ref},F_{e}).$ (19)
-We note that the second-last equation ($SE(3)$ left-invariance of the elastic wrench) comes from Lemma 1 of [11]. Finally, from the result of Proposition 2 of [11], it follows that
-$h^{s}_{\theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})=\text{Ad}_{g_{l}^{\text{-}1}}^{T}h_{\theta}^{s}(g,g_{ref},F_{e}),$ (20)
-where Ad is an Adjoint operator, $\text{Ad}_{g_{l}^{\text{-}1}}^{T}$ is a representation for $se^{*}(3)$ (wrench) domain for the group action $g_{l}$ [11], and superscript $s$ denotes a vector represented in the spatial frame.
-∎
+Suppose that the Assumption 1 and 2 hold. The EquiContact Policy $h_{\theta}$ is equivariant if it is described relative to the spatial frame.
+The proof is shown in the Appendix. -B.
+Algorithm 1 Inference Procedure of EquiContact
+1:Diff-EDF $f_{\theta_{1}}$, G-CompACT $\pi_{\theta_{2}}$, Task $\in$ {pick, place}
+2:Get scene and grasp point cloud $\mathcal{O}^{scene}$, $\mathcal{O}^{grasp}$
+3:Run Diff-EDF for reference frame $g_{{}_{EDF}}\!=\!f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$
+4:Move the end-effector near the reference frame and
+5:   initialize EquiContact $\pi_{\theta_{2}}$
+6:for each inference timestep $k$ do
+7:  Get current sensor values $g(k)$, $F_{e}(k)$, $I_{w}(k)$
+8:  Calculate GCEV $e_{{}_{G}}(k)=e_{{}_{G}}(g(k),g_{{}_{EDF}})$ (4)
+9:  Run G-CompACT:
+10:      $(g_{rel},K_{p},K_{R})(k)=\pi_{\theta_{2}}(e_{{}_{G}},F_{e},I_{w})(k)$ (3)
+11:  Calculate desired pose $g_{d}(k)=g(k)g_{rel}(k)$
+12:  Update $(g_{d},K_{p},K_{R})(k)$ for GAC loop
+13:  Run GAC realizing desired dynamics (9), (10)
+14:end for
 IV-D Extensions to Pick Tasks
@@
 V-A Benchmark Results
-To validate the effectiveness of our proposed EquiContact framework and its underlying design principles, we compare EquiContact against three benchmark approaches: ACT with world-frame observation and action representation, executed with and without GAC, and CompACT [17].
-We note that the GAC serves as a standard admittance control when the platform’s pose is fixed, i.e., untilted.
-The observation and action representation of the benchmarks and the proposed EquiContact are listed in Table I.
-We conducted a series of benchmark tests to demonstrate the effectiveness of each component in the proposed approach.
-The overall results are summarized in Table I.
+To evaluate the effectiveness of EquiContact and its core design principles, we compare it against three baselines: ACT with world-frame observations and actions, executed with and without GAC, and CompACT [17].
+Note that GAC reduces to standard admittance control when the platform is fixed, i.e., not tilted.
+Table.
+I summarizes the observation/action representations used in each method and reports the benchmark results across all setups.
 TABLE I: Success rates of the insertion policies in real-world experiments for the proposed and benchmark approaches.
@@
 V-A1 Demonstration of Compliance
+As the importance of left-invariant compliant control action has already been verified in [11], we focus on verifying the necessity of compliant control action.
 We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC).
@@
 Due to the safety issue, we limited the number of trials without GAC to 10.
-This result demonstrates that the presence of compliant control is almost a decision factor between success and failure for contact-rich tasks.
+This result demonstrates that compliant control is nearly a deciding factor between success and failure in contact-rich tasks.
 To further analyze the effects of compliant gain control, we compare ACT with GAC and CompACT under in-distribution (In-Dist.) platform settings.
@@
 This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance.
-The result of applying CompACT to the translationally unseen cases is shown in the Table I $3$rd row - Flat Platform (OOD).
+The result of applying CompACT to the translationally unseen cases is shown in the Table I 33rd row - Flat Platform (OOD).
 Note that the flat platform is only utilized, meaning it is only randomized translationally.
@@
 However, as the whole task is formulated in sequential stages, the error in the previous stage tends to propagate to the following stages, leading to slightly increased failure cases.
-From these experimental results, we draw the following key takeaway.
-Remark 1 (Anchoring a Localized Policy on the Reference Frame for Equivariance).
-The G-CompACT policy $\pi_{\theta}$ operates exclusively on localized inputs: GCEV, FT values in the end-effector frame, and images from wrist cameras.
-Its outputs, relative poses, and admittance gains are likewise defined in the local end-effector frame.
-The localized policy is anchored to a reference frame generated by the Diff-EDF planner.
-This architectural design induces spatial generalization and $SE(3)$-equivariance.
-Under this principle, one may hypothesize the reason for the spatial generalizability of recent vision-language-action models, which utilize both wrist and external cameras.
-The wrist camera provides a localized policy, anchored in the reference frame obtained from the external cameras and proprioceptive information.
-In fact, learning a reference frame from proprioceptive information (e.g., joint position) and the external camera is subtle because of its black-box nature.
-In contrast, our method, which utilizes a point cloud-based reference frame approach, is more interpretable.
-Thus, we propose “anchoring localized policies on globally estimated reference frames” as a general framework for the divide-and-conquer philosophy to enhance spatial generalization and interpretability.
-In our case, this structure also enables provable $SE(3)$ equivariance via differential geometric design.
-V-B Failure Cases and Performance Recovery via Data Augmentation
+V-B Limitations and Discussions
 TABLE III: Success Rates of the G-CompACT (insertion) trained with the base dataset and augmented dataset for the flat platform under the presence of visual distractor scenario and tilted platform with large angle cases.
@@
 Tilted Platform ($45^{\circ}$) 5 / 10 9 / 10
-Although the EquiContact framework is vision-to-force equivariant in theory, this guarantee holds only under Assumptions 1 and 2.
-In particular, Assumption 2, which requires left-invariance of the visual features, is more challenging to enforce in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias.
-As a result, the proposed G-CompACT algorithm shows degraded performance in scenarios with visual distractors or on a more severely tilted platform $(45^{\circ})$ – see Table III (“Base Dataset” column).
-Note that the larger tilting angle results in more out-of-distribution images, as the wrist cameras begin to capture more unseen background scenes, as illustrated in Fig.
-3.
-To accommodate this, we augment the base dataset with $20$ demos collected using the visual distractor on a fixed, flat platform and $20$ demos collected from a $30^{\circ}$ angle of the tilted platform.
-After training with the augmented dataset, the success rates return to normal levels, as shown in the “Aug.
-Dataset” column of Table III, demonstrating the performance recovery achieved through data augmentation.
+V-B1 Limitation on Left-invariant Visual Features
+Although EquiContact is theoretically vision-to-force equivariant, this guarantee holds only under Assumptions 1 and 2.
+In particular, Assumption 2, which requires left-invariant visual features, is challenging to ensure in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias.
+Consequently, G-CompACT shows degraded performance in settings with visual distractors or severe tilting ($45^{\circ}$), as shown in Table III (“Base Dataset”).
+The degradation arises because larger tilt angles expose unseen background regions in the wrist camera view (see Fig.
+3), resulting in out-of-distribution observations.
+This can be mitigated by augmenting the training data with visual distractors and demonstrations collected from a tilted angle.
+To demonstrate this, we added $20$ demos with distractors on a flat platform and 20 demos from a $30^{\circ}$ tilted setup.
+As shown in the “Aug.
+Dataset” column of Table III, the success rates have recovered to original levels.
+Enforcing left-invariance on RGB images via architectural design remains an open challenge.
+Recent works have explored inducing 3D equivariance from 2D images [32, 33].
+Another promising direction involves using pretrained foundation models, such as CLIP and DINO, which may exhibit implicit equivariance due to large-scale and diverse training.
+Language-guided features (e.g., using the prompt “peg”) can help extract relevant features across views.
+Alternatively, real-time segmentation models can be used to focus only on the object of interest.
+We will investigate enforcing left-invariance of images in future work.
+V-B2 Hardware and Algorithmic Limitation
+A key challenge in implementing equivariant policies for robotic manipulation is handling singularities [10].
+When the robot is near singular, pose tracking accuracy degrades because controllers sacrifice tracking to avoid singularity, leading to poorly executed policy commands.
+This results in a distributional shift, which is a critical limitation for imitation learning.
+From our experimental demonstration perspective, a small workspace of the Diff-EDF (Fig.
+2) was a drawback due to its limited receptive field.
+A combination of the singularity, joint limits, and small workspace resulted in limited demonstration of our approach, not covering a complete $SE(3)$ transformation.
+V-C Validating Feasibility to Other Contact-rich Tasks
+To further validate the EquiContact framework, we test it on two additional contact-rich tasks: screwing and surface wiping (see Fig.
+5).
+In the screwing task, the robot aligns and screws a peg; in the wiping task, it erases a line from a whiteboard.
+For simplicity, we use a white background to eliminate visual distractions and assume that reference frames are known.
+The reference frame for screwing is the end-effector pose at full insertion; for wiping, it is the center of the board.
+As in the PiH setup, demonstrations are collected on a fixed platform and evaluated on out-of-distribution configurations, including tilted platforms.
+We collected 50 demonstrations per task.
+Results in Table IV show consistent success rates across all conditions, confirming EquiContact’s spatial generalizability.
+Figure 5: (Left) Screwing task, (Right) Surface wiping (erasing) task. The same platform structure of the PiH is used.
+TABLE IV: Success Rates of the G-CompACT for screwing and surface wiping tasks. The evaluation is conducted with the ground-truth reference frames. Both tasks are trained with $50$ demos.
+Test Scenario Screwing Surface Wiping
+Flat Platform (OOD) 10 / 10 10 / 10
+Tilted Platform ($30^{\circ}$, OOD) 9 / 10 9 / 10
 VI Conclusion
@@
 By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations.
-We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations and validated its effectiveness through real-world experiments on PiH benchmarks.
+We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations, validated its effectiveness through real-world experiments on PiH benchmarks and its feasibility towards screwing and surface wiping tasks.
 Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates.
@@
 Supplementary Material
-SI Implementation Details
+Implementation Details
 In this section, we provide the details of the hardware code implementation.
 Since every hardware system has a different implementation, our goal is to provide a comprehensive overview and a guide to our code distribution.
-We will use the exact names of the nodes to enable readers to easily understand our codebase.
-SI-A Implementation Pipeline
+We will use the exact names of the nodes to enable readers to understand our codebase easily.
+The hardware experimental implementation code is provided in https://github.com/Joohwan-Seo/EquiContact-Experiment, and the simulation implementation code in https://github.com/Joohwan-Seo/EquiContact-Simulation.
+S-A Implementation Pipeline
 The pipeline visualized in Fig.
@@
 Figure S1: Whole pipeline implemented (via client named as FullPipelineClient) with ROS2 is presented.
-SI-B Diffusion-EDF
-SI-B1 Implementation Heuristics
+S-B Diffusion-EDF
+S-B1 Implementation Heuristics
 Instead of following the original Diff-EDF pipeline, which utilizes pick-and-place models, we used two pick models.
@@
 where $\bar{R}_{place}$ is the averaged rotation of $R_{place}^{i}$, and $\bar{p}_{tip,place}$ is the averaged position of the tip calculated similarly from the picking case.
-SI-B2 Accuracy of the Diff-EDF
-Here, we present detailed error values of the Diff-EDF evaluated on the training dataset.
-We note that we followed the implementation pipeline of the original Diff-EDF package, but just utilized the two pick models as described in Section SI-B1.
-The resulting root mean squared error (RMSE) values are shown in Table S1.
-The rotational errors are calculated via $R_{e}=R_{gt}R_{pick}^{T}$ and converted to $xyz$ euler angle representation, where $pick$ is replaced by $place$ for the placing task.
-TABLE S1: RMSE error values of Diff-EDFs. The dimension of translational error is $\mathrm{mm}$ and the rotational error is $\mathrm{deg}$.
-Pick Place
-Translation $x$ ($\mathrm{mm}$) 16.76 4.981
-Translation $y$ ($\mathrm{mm}$) 5.886 7.422
-Translation $z$ ($\mathrm{mm}$) 10.79 5.236
-Rotation $x$ ($\mathrm{deg}$) 19.88 13.74
-Rotation $y$ ($\mathrm{deg}$) 40.55 18.99
-Rotation $z$ ($\mathrm{deg}$) 106.7 91.41
-As mentioned in Section I, the error in translation ranges from $5\mathrm{mm}$ to $15\mathrm{mm}$, and rotational error can range up to $40^{\circ}$.
-We note that the $z$-directional rotation error can be arbitrarily high in our case because our PiH task is symmetric.
-SI-C Tips on applying Geometric Admittance Control
+S-C Tips on applying Geometric Admittance Control
 Here, we present a few tips on the application of geometric admittance control. The GAC is implemented as a part of a ROS2 node, running with $200\mathrm{Hz}$.
-SI-C1 ROS2 Implementation
+S-C1 ROS2 Implementation
 The Robot Node acts as the central controller, receiving and executing desired pose commands using a Geometric Admittance Control (GAC).
@@
 Within its main loop, the controller sends the intended pose and processed FT sensor data to the GAC function, which generates the compliant commands for the robot’s low-level motion controller.
-SI-C2 Handling Force-Torque Sensor Data
+S-C2 Handling Force-Torque Sensor Data
 The first consideration is how to handle the Force-Torque (FT) sensor attached to the end-effector.
@@
 We designed a simple 1st-order low-pass filter and implemented it using a state-space representation in discrete time.
-SI-C3 Implementation Tips
+S-C3 Implementation Tips
 As mentioned in Section.
@@
 If a torque controller is readily available, then directly implementing a geometric impedance controller (GIC) is recommended, but joint friction compensation needs to be handled appropriately.
-The end-effector velocity control can be straightforwardly applied from (IV-B1), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command.
+Given the desired dynamics (9), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
+$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
+$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (16)
+where $T_{s}$ is a sampling time ($5\mathrm{ms}$ in our case for GAC) and $\hat{(\cdot)}$ denotes a hat-map.
+The end-effector velocity control can be straightforwardly applied from (S-C3), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command.
 One caveat here is the $V^{b}$, which is body-frame end-effector velocity.
@@
 For Python, the Jacobian matrix function can also be wrapped into executable .so file using Matlab code builder and executed in the loop for fast computation (well below $1\mathrm{ms}$).
-SI-D Teleoperation Details
+S-D Teleoperation Details
 The architecture of our teleoperation system is explained to make it easier to reproduce and adapt our work.
@@
 This allows greater adaptation to different robotic platforms and control algorithms.
-Two ROS nodes make up the majority of the architecture: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
+The teleoperation function is implemented by two ROS nodes: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
 The teleoperation node acts as the principal interface between the human operator and the robot controller.
@@
 These gain profiles are published as action messages, together with poses and gripper states.
-SI-E G-CompACT Training Details
+S-E G-CompACT Training Details
 During the training of the G-CompACT, we added the noise to the known fixed pose of the hole and peg.
@@
 where $n_{p}\sim\mathcal{U}(-0.02,0.02)\in\mathbb{R}^{3}$ and $n_{r}\sim\mathcal{U}(-8^{\circ},8^{\circ})\in\mathbb{R}^{3}$, which denotes $2\mathrm{cm}$ of translational and $8^{\circ}$ (degree) of rotational error.
+The impact of adding noise to the reference frame is further shown in Appendix.
+-C.
 To encode the element on $SE(3)$ for learning, we used position & rotation vector representation for G-CompACT.

### v3

@@
 This paper presents a framework for learning vision-based robotic policies for contact-rich manipulation tasks that generalize spatially across task configurations.
-We focus on achieving robust spatial generalization of the policy for the peg-in-hole (PiH) task trained from a small number of demonstrations.
+We focus on achieving robust spatial generalization of the policy for the contact-rich tasks trained from a small number of demonstrations.
 We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant Action Chunking Transformers, G-CompACT).
@@
 We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance.
-Real-world experiments on PiH, screwing, and surface wiping tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles.
-The experimental videos can be found on the project website: https://sites.google.com/berkeley.edu/equicontact
+Real-world experiments on peg-in-hole (PiH), screwing, and surface wiping tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles.
+The experimental videos will be attached as supplementary material, and the codes will be released.
 I Introduction
@@
 An alternative line of recent research focuses on leveraging symmetry—particularly equivariance—to enhance spatial generalizability, thereby improving sample efficiency during training [22, 19].
-This approach requires fewer data but comes with its own challenges.
+This approach requires less data but comes with its own challenges.
 Equivariant neural networks, being of a more specialized nature, are often not as well-developed and are more computationally intensive than their non-equivariant counterparts, making real-time and large-scale deployment more difficult.
@@
 Since both the high-level planner and low-level variable impedance controller are equivariant, they can be combined to form a vision-to-force equivariant policy.
-However, in practice, Diff-EDF’s placement accuracy proved insufficient for tight insertion tasks, which require sub-millimeter precision (details in Appendix -A).
+However, in practice, Diff-EDF’s placement accuracy proved insufficient for precision tasks, which require sub-millimeter precision (details provided in Appendix AIII-A.
 This revealed a key limitation: high-level vision planners may capture global structures but struggle with precision and contact-sensitive execution.
-From this insight, we introduce an intermediate layer between the planner and the low-level controller, which provides real-time visual feedback to correct the residual errors of the high-level planner.
+Henceforth, we introduce an intermediate layer between the planner and the low-level controller, which provides real-time visual feedback to correct the residual errors of the high-level planner.
 In this paper, we propose EquiContact, a hierarchical $SE(3)$ vision-to-force equivariant policy for spatially generalizable, contact-rich tasks.
@@
   2. 2.
-    We identify three key principles for spatially generalizable contact-rich manipulation: left-invariant compliant control action (via GAC [11]), localized policy (left invariance), and induced equivariance.
+    We identify three key principles for spatially generalizable contact-rich manipulation: (1) left-invariant compliant control action (via GAC [19]), (2) localized policy (left invariance), and (3) induced equivariance.
 These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [22].
@@
     We demonstrate that EquiContact achieves near-perfect success rates and spatial generalizability when these conditions are met in real robot experiments involving peg-in-hole, screwing, and surface wiping tasks.
-From these key principles, we propose a general framework to enhance the spatial generalization and interpretability of vision-based policies, namely, “anchoring localized policy on globally estimated reference frame.” We emphasize that our work provides complementary insights to recent trends of robot learning [1, 8, 16, 2] that aim to build generalist policies from a large-scale demo dataset.
-Our principles offer structural guidelines to improve spatial generalizability through $SE(3)$ equivariance.
+From these key principles, we propose a general framework to enhance the spatial generalization and interpretability of vision-based policies, namely, “anchoring localized policy on globally estimated reference frame.” We emphasize that our work provides complementary insights to recent trends in robot learning [13, 9, 5, 1] that aim to build generalist policies from large-scale demonstration datasets.
+Our principles provide structural guidelines for improving spatial generalizability via $SE(3)$ equivariance.
 II Related Works
@@
 Particularly, there are two dominant methods for visuomotor servoing: Action Chunking with Transformers (ACT) [37] and Diffusion Policy (DP) [4].
-ACT utilizes the conditional variational autoencoder (CVAE) method as a generative model, whereas DP employs denoising diffusion as its generative model.
+ACT uses a conditional variational autoencoder (CVAE) as its generative model, whereas DP uses denoising diffusion.
 ACT and DP have been extended to other approaches, including compliance and force-reactive behaviors [8, 7, 32], as well as structural improvements [14, 5, 29].
@@
 Importantly, compared to [3, 17], we explicitly link the choice of reference frame to the equivariance property, and unlike [36], which only handles translational transformations, our method can deal with full $SE(3)$ transformations of the reference frame.
-Figure 1: We propose an EquiContact, a hierarchical, provably $SE(3)$ vision-to-force equivariant policy for spatially generalizable contact-rich tasks.
-The proposed EquiContact consists of G-CompACT (Section.
-IV-B) and Diffusion-EDF (Section.
-IV-A).
-The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen task transformations during evaluation.
-The overall EquiContact algorithm is summarized in Algorithm 1.
 III Problem Definition
+Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented.
+22 external cameras with calibrated extrinsics and 22 wrists cameras are installed.
+The workspace shown is the Diff-EDF workspace.
+(Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance.
+(Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
 In this paper, we aim to identify the key structural components required for learning policies that generalize spatially in contact-rich manipulation tasks.
@@
 These principles are validated through specific data collection and evaluation setups, as detailed below.
-Unlike prior work [11] that assumes a known hole pose and a pre-grasped peg, we consider a more general setup: the robot must first grasp the peg and then perform insertion using only vision and proprioception, as illustrated in Fig.
+Unlike prior work [19] that assumes a known hole pose and a pre-grasped peg, we consider a more general setup: the robot must first grasp the peg and then perform insertion using vision, proprioception, and task description in text, as illustrated in Fig.
 2.
-We assume only that the peg is upright in the scene.
-Inspired by the idea that many manipulation tasks can be viewed as sequences of pick-and-place actions, we formulate insertion as a compliant pick-and-place task.
-However, high-level vision planners often lack the precision needed for tight-tolerance tasks like PiH, which in our case requires $1\mathrm{mm}$ accuracy.
+We assume the peg is upright and that the hole’s yaw angle is known within $90^{\circ}$ range.
+Given an initial estimate, we resolve the orientation by selecting the closest angle among the four symmetric candidates (e.g., $\psi,\psi+90^{\circ},\psi+180^{\circ},\psi+270^{\circ}$).
+As mentioned earlier, high-level vision planners often lack the precision needed for tight-tolerance tasks like PiH, which in our case requires $1\mathrm{mm}$ accuracy.
 To accommodate this, we propose a low-level compliant policy that:
@@
 To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location, where its objective is to train a policy that performs nearly perfectly in the trained scenario.
-We then evaluate benchmark and proposed methods—trained solely on these limited demonstrations—across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
-Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented.
-22 external cameras with calibrated extrinsics and 22 wrists cameras are installed.
-(Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance.
-(Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
+We then evaluate benchmark and proposed methods, trained solely on these limited demonstrations, across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
+Further, we demonstrate that our proposed approach can adapt to extreme task transformations as shown in Fig.
+3.
+Finally, the validity of the proposed EquiContact method is shown in other contact-rich tasks, such as the screwing task and surface-wiping tasks.
+Figure 3: Extreme task transformations.
+(Left) $90^{\circ}$ transformation in $x$ axis.
+(Middle) $135^{\circ}$ transformation in $x$ axis.
+(Right) $45^{\circ}$ transformation in $y$ axis, facing towards the camera.
 IV Solution Approach
-We now describe our proposed EquiContact policy, which integrates a high-level vision planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant ACT (G-CompACT).
-G-CompACT consists of a behavior cloning module based on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC).
-As discussed in Section III, we focus on the insertion task, with extension to picking addressed later in the paper.
-The Diff-EDF and G-CompACT are trained separately and combined at inference, as shown in Fig.
-1.
-Conceptually, EquiContact follows a simple yet powerful principle: “anchoring a localized policy on a globally estimated reference frame.” In our framework, G-CompACT serves as the localized low-level policy, operating solely on observations defined in the current end-effector frame; wrist camera RGB images, force-torque (FT) signals, and a geometrically consistent error vector (GCEV) $e_{{}_{G}}$.
-It outputs the desired relative pose and admittance gains, also in the end-effector frame.
-As both inputs and outputs are defined locally, G-CompACT functions as a fully localized policy.
-The high-level planner, Diff-EDF, estimates the pose of the target (e.g., hole) in the world frame.
+We introduce the EquiContact framework in this Section.
+The EquiContact framework integrates a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) with a low-level compliant visuomotor policy (Geometric Compliant ACT, G-CompACT) and geometric admittance control (GAC) at the lowest level.
+The Diff-EDF gets the point cloud inputs from external cameras to generate reference frames.
+Based on the estimated reference frames, the G-CompACT process the real-time proprioceptive and wrist camera feedback to output desired poses and admittance gains.
+In what follows, the GAC module outputs the geometrically consistent compliant motion from desired poses and admittance gains to enable equivariant force interaction.
+We first focus on the insertion task, with extension to picking addressed later in the paper.
+Conceptually, EquiContact follows a simple yet powerful principle: “anchoring a localized policy on a globally estimated reference frame.” In our framework, G-CompACT serves as a fully localized low-level policy, operating solely on observations defined in the current end-effector frame, and actions defined in the end-effector frame – See Fig.
+4.
+The high-level planner, Diff-EDF, estimates the pose of the target (e.g., hole) in the world (global) frame.
 At the inference, the robot moves near the estimated reference frame, and G-CompACT is activated to perform compliant motion using only local feedback.
@@
 This divide-and-conquer design provides a general framework for enhancing both spatial generalization and policy interpretability for contact-rich, and more broadly, general manipulation tasks.
-In the remainder of this section, we formalize this spatial generalization property as $SE(3)$ equivariance and show how EquiContact satisfies this by construction through a rigorous mathematical derivation.
-We note that $\theta$, $\theta_{i}$, and $\phi$ with $i=1,2,3$ denote a learnable parameter of the neural network, without loss of generality.
-We also note that we follow the notation and definition conventions of [11, 30].
-IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
-Diff-EDF [14] is a diffusion-based manipulation policy with a bi-equivariant structure on $SE(3)$ transformations for pick-and-place tasks.
-Diff-EDF takes two point clouds as input: the scene point cloud $\mathcal{O}^{scene}$ captured by two external RGBD cameras with calibrated extrinsics (Fig.
-2), and the gripper point cloud $\mathcal{O}^{grasp}$ in the end-effector frame, with or without a grasped object.
-Diff-EDF outputs a pose $g_{{}_{EDF}}\in SE(3)$ representing the estimated target pose of the object of interest, which in our case is the hole position and is defined as:
-$g_{{}_{EDF}}=f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$ (1)
-where the neural network $f_{\theta_{1}}$ implements the mapping.
-Due to the left-equivariant property, it follows that: $g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
-Let $\mathcal{O}^{ref}\subset\mathcal{O}^{scene}$ denote the subset of points corresponding to the object of interest (e.g., the hole structure, see also Fig.
-3).
-The model is designed to satisfy the following left-equivariance property on the point cloud characterization of the target object of interest:
-Assumption 1 (Left-Equivariance of Diff-EDF).
-$\begin{split}g_{{}_{EDF}}&=f_{\theta_{1}}(\mathcal{O}^{ref},\mathcal{O}^{grasp})\\ &\implies g_{l}\cdot g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{ref},\mathcal{O}^{grasp}),\end{split}$ (2)
-To meet this assumption, we randomize the pose of the hole-platform assembly during training, both translationally and rotationally, with visual distractors in the background.
-This encourages the model to focus on $\mathcal{O}^{ref}$, not the rest of the objects in the scene.
-While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and has high robustness to left $SE(3)$ transformations.
-To train Diff-EDF, the scene and grasp point clouds are collected together with the target reference frames, which represent the desired poses of the end-effector for pick-and-place operations.
-$15$ demonstrations were collected for the Diff-EDF: $12$ samples of the flat platform and $3$ samples of the tilted platform, both translationally and rotationally randomized.
-The training process of Diff-EDF follows the procedure in [14].
-Note that arbitrary peg transformations relative to the gripper can be handled via the right-equivariance feature of the Diff-EDF; however, due to the complications mentioned in Section.
-III, we enforced a consistent grasp and only utilized the left-invariance property.
-We refer the readers to [14, 10] for full details of Diff-EDF.
-IV-B Geometric Compliant control Action Chunking with Transformers (G-CompACT)
-IV-B1 G-CompACT
+In the remainder of this section, we formalize this spatial generalization property as $SE(3)$ equivariance and show how EquiContact satisfies this by its design choices.
+IV-A Geometric Compliant control Action Chunking with Transformers (G-CompACT)
 G-CompACT is based on the Action Chunking with Transformer (ACT), which is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks [37].
-To make G-CompACT spatially equivariant, we follow the principles proposed in [11]: left-invariant policy and policy representation in the end-effector body frame.
-To achieve this, we structure the observations and actions of the G-CompACT as described below.
-First, the observations are given by, (1) Geometrically Consistent Error Vector (GCEV) $e_{{}_{G}}$ proposed in [11], (2) FT sensor in the end-effector frame $F_{e}$ to capture contact behaviors, and (3) RGB images $I_{w}=\{I_{w,1},I_{w,2}\}$ from wrist cameras (see Fig.
-3).
+To make G-CompACT spatially equivariant, we follow the principles proposed in [19]:
+  • Left-invariant policy, and
+  • Policy representation in the end-effector body frame
+We have designed the G-CompACT architecture to achieve this properties as described in this chapter. The overall structure of the G-CompACT is also summarized in Fig. 4.
+Figure 4: G-CompACT architecture is presented.
+The G-CompACT $\pi_{\theta}$ receives GCEV $e_{{}_{G}}$ (2) and F/T sensor value $F_{e}$ as proprioceptive inputs, along with two wrist camera images $I_{w}$.
+The wrist camera images are fed to the CLIP-ResNet50 visual backbone, followed by the Feature-wise Linear Modulation (FiLM) layer.
+The FiLM layer is modulated by the text tokens $t_{text}$ from the (frozen) CLIP text encoder, which processes the task descriptions.
+The proprioceptive inputs and the latent features of modulated vision $z$ are then fed to the transformer decoder $D_{\psi}$, which outputs the action signals $a$.
+Note that we omitted the style variable for the transformer decoder, and 0 values are used during inference.
+The observations are given by, (1) Geometrically Consistent Error Vector (GCEV) $e_{{}_{G}}$ proposed in [19], (2) FT sensor in the end-effector frame $F_{e}$ to capture contact behaviors, and (3) RGB images $I_{w}=\{I_{w,1},I_{w,2}\}$ from wrist cameras (see Fig.
+5).
 For actions, we choose (1) relative pose from the current end-effector frame $g_{rel}$, and (2) admittance gains for Geometric Admittance Control (GAC) $(K_{p},K_{R})$.
 The details of GAC and the definition of gains will be provided later in the Section.
-Formally, let the G-CompACT method be $\pi_{\theta_{2}}$ such that
-$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq(g_{rel},K_{p},K_{R})(k),\quad o(k)\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (3)
+Formally, the G-CompACT method $\pi_{\theta}$ can be written as:
+$\begin{split}a(k)&=\pi_{\theta}(o(k)),\quad\text{where}\\ a(k)&\triangleq(g_{rel},K_{p},K_{R})(k),\quad o(k)\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (1)
 where $a(k)$ denotes the actions, and $o(k)$ denotes the observation at time step $k$.
-In fact, our G-CompACT also outputs the actions of chunk size $N$ following the standard action-chunking strategy.
-However, for the compact presentation, we will only consider the single-step action after undergoing proper processing, such as a temporal ensemble.
-The GCEV $e_{{}_{G}}(g,g_{{}_{EDF}})$ is defined as
-$e_{{}_{G}}(g,g_{{}_{EDF}})=\begin{bmatrix}R^{T}(p-p_{{}_{EDF}})\\ (R_{{}_{EDF}}^{T}R-R^{T}R_{{}_{EDF}})^{\vee}\end{bmatrix}$, (4)
-where $g=(p,R)\in SE(3)$ is a current end-effector pose, $g_{{}_{EDF}}=(p_{{}_{EDF}},R_{{}_{EDF}})$ is the reference frame predicted by Diff-EDF, and $(\cdot)^{\vee}$ denotes the vee-map, a mapping from $so(3)$ (Lie algebra of $SO(3)$) to $\mathbb{R}^{3}$.
-The physical meaning of GCEV is an error vector in the current end-effector frame between the current end-effector pose and the reference frame, considering the $SE(3)$ manifold structure, i.e., non-Euclidean.
+Although the G-CompACT outputs the actions of chunk size $N$, we will only consider the single-step action after proper processing, such as a temporal ensemble, for notational compactness.
+The GCEV $e_{{}_{G}}(g,g_{ref})\in\mathbb{R}^{6}$ is defined as
+$e_{{}_{G}}(g,g_{ref})=\begin{bmatrix}R^{T}(p-p_{ref})\\ (R_{ref}^{T}R-R^{T}R_{ref})^{\vee}\end{bmatrix}$, (2)
+where $g=(p,R)\in SE(3)$ is a current end-effector pose, $g_{ref}=(p_{ref},R_{ref})\in SE(3)$ is the reference frame estimated by the global estimator, e.g., Diff-EDF, and $(\cdot)^{\vee}$ denotes the vee-map, a mapping from $so(3)$ (Lie algebra of $SO(3)$) to $\mathbb{R}^{3}$.
+The physical meaning of GCEV is an error vector between the current end-effector frame and the reference frame, defined on the current end-effector frame.
+As will be elaborated in Appendix AI, the proprioceptive signals $e_{{}_{G}}$ and $F_{e}$ are left-invariant.
 For the details of GCEV $e_{{}_{G}}$, we refer to [19, 20].
-Since the images $I_{w}$ are fed to the transformer decoder after being passed to the visual encoder structure, one can further consider the G-CompACT as
-$\begin{split}a(k)&=\pi^{\prime}_{\theta_{3}}(e_{{}_{G}},F_{e},z)(k),\quad\text{where}\;z(k)=\mu_{\phi}(I_{w}).\end{split}$ (5)
-where $\mu_{\phi}$ denotes the visual encoder, e.g., ResNet, and $z$ is the latent variable from the visual encoder.
-We now introduce the following assumption regarding the left-invariant visual representation.
-Assumption 2 (Left-invariant Visual Features).
-The features from visual encoder $\mu_{\phi}$ are left invariant, i.e.,
-$z(k)=\mu_{\phi}(g_{l}\circ I_{w})=\mu_{\phi}(I_{w}),\;\;\forall g_{l}\in SE(3).$ (6)
-Note that we refer to $g_{l}$ as a left-group action [10].
-The left-group action applied on the wrist-camera images is illustrated in Fig.
-3.
+The images $I_{w}$ are fed to the transformer decoder $\mathcal{D}_{\psi}$ after being processed by the vision encoder structure $\mu_{\phi}$; therefore, one can further represent G-CompACT as
+$a(k)=\mathcal{D}_{\psi}(e_{{}_{G}},F_{e},z)(k)=\mathcal{D}_{\psi}(e_{{}_{G}},F_{e},\mu_{\phi}(I_{w}))(k),$ (3)
+where $z=\mu_{\phi}(I_{w})$ is a visual feature from the vision encoder.
+To satisfy the left-invariant condition of G-CompACT, the features from the vision encoder $z$ need to be invariant to the left-transformation of the image.
+We formalize the left-invariant visual feature condition as the following assumption.
+Assumption 1 (Approximately Left-invariant Visual Features).
+The visual encoder $\mu_{\phi}$ produces features that are approximately left-invariant to task transformations, i.e.,
+$\mu_{\phi}(g_{l}\circ I_{w})\approx\mu_{\phi}(I_{w}),$ (4)
+$\forall g_{l}\in SE(3)$ that preserves local task geometry.
+Here, $\approx$ notation denotes invariance up to a bounded representation error that does not affect the policy’s qualitative behavior.
+Note that we refer to $\circ$ as a group action [22].
+The left-group action applied to the wrist-camera images $g_{l}\circ I_{w}$ is illustrated in Fig.
+5.
 The meaning of the visual representation $z$ being left-invariant is that the vision encoder $\mu_{\phi}$ is trained to focus only on group action invariant features, such as the flat surface surrounding the hole on the platform.
-The satisfaction of Assumption 2 can be challenging.
-To satisfy this assumption, we have designed our platform and the surface surrounding the hole assembly so that there are sufficient surface features and the cameras primarily see the surface surrounding the hole, not the lower parts of the platform.
-This engineering choice was somewhat ad hoc, as there is no guarantee or sufficient inductive bias to encourage the desired behavior.
-We will demonstrate in later experimental results that this assumption may not apply in certain cases.
+To satisfy this assumption, we use language grounding to extract vision features that are correlated with the language description.
+The core insight behind this is that language tokens encode object identity rather than pose, and thus provide a conditioning signal that is invariant to global $SE(3)$ transformations of the scene.
+For example, a “peg” is still “peg” no matter from which view it is seen.
+Specifically, the pretrained CLIP-ResNet5050 (CLIP-RN50) [16] is utilized for the vision encoder, and the CLIP text encoder is also employed.
+Although the pretrained CLIP-RN50 was used, it was fully retrained (details provided in Appendix AII); thereby, it served as a good initialization.
+We provide the task descriptions to the text encoder to obtain text tokens $t_{text}$.
+The vision feature $z_{raw}$ of CLIP-RN50 backbone is then modulated using feature-wise linear modulation (FiLM, [15]) layer from $t_{text}$ via
+$z=\beta(t_{text})z_{raw}+\gamma(t_{text}),$ (5)
+where $\beta$ and $\gamma$ are trainable FiLM layer.
+Using FiLM, the vision features are suppressed or highlighted to align with task-relevant semantic concepts, empirically encouraging approximate left-invariance with respect to workspace transformations.
 Figure 5: Effects of the left group action $g_{l}$ to the end-effector pose $g$ and the reference frame $g_{ref}$, and to the wrists cameras $I_{w,1}$ and $I_{w,2}$.
-As the left group action is applied on the end-effector, the wrist cameras start to see backgrounds other than the optical table.
-$\{s\}$ denotes a spatial frame, i.e., robot base frame or world frame.
-The following proposition shows the left-invariance of the G-CompACT method in the end-effector frame.
-Proposition 1 (Left-invariance of G-CompACT).
-Suppose that the Assumption 2 holds. Then,
-$a(k)=\pi_{\theta_{2}}(g_{l}\circ o(k))=\pi_{\theta_{2}}(o(k)).$ (7)
-The proof is presented in Appendix. -B. In what follows, we present an equivariant property of the pose signal produced by G-CompACT when described in the spatial frame.
+As the left group action is applied to the end-effector and the target object, the wrist cameras start to see the backgrounds of arbitrary lab objects.
+Under the satisfaction of Assumption 1, i.e., ideal left-invariant vision feature condition, the following proposition shows the left-invariance of the G-CompACT method in the end-effector frame.
+Proposition 1 (Left-invariance of G-CompACT under ideal invariant visual features).
+Suppose Assumption 1 holds. Then the G-CompACT policy $\pi_{\theta}$ is left-invariant in the end-effector frame, i.e.,
+$\pi_{\theta}(g_{l}\circ o(k))=\pi_{\theta}(o(k)),\quad\forall g_{l}\in SE(3).$ (6)
+The proof is presented in Appendix AI. A remark is provided.
+Remark 1 (Implication of Proposition 1).
+Proposition 1 highlights a sufficient structural condition for spatial generalization: if the visual encoder and proprioceptive signals can be made (approximately) left-invariant to $SE(3)$ task transformations, then the resulting closed-loop policy inherits left-invariance by construction.
+This motivates learning or enforcing representations that satisfy Assumption 1.
+In what follows, we present an equivariant property of the pose signal produced by G-CompACT when described in the spatial frame.
 Corollary 1 ($SE(3)$ Equivariance of G-CompACT).
-The G-CompACT $\pi_{\theta_{2}}$ represented in the spatial frame satisfies the following equivariance property:
-$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\pi_{\theta_{2}}(g_{l}\circ o(k)).\end{split}$ (8)
-Proof.
-Notice that the desired pose signal in the spatial frame $g_{d}$ is obtained via $g_{d}=gg_{rel}=g\pi_{\theta_{2}}(o(k))$.
-Then, utilizing the left-invariance property, $g_{l}g\cdot\pi_{\theta_{2}}(g_{l}\circ o(k))=g_{l}g_{d}$.
-∎
-To train a G-CompACT, we collect the expert demonstration using teleoperation on the fixed pose of the platform.
-During data collection, the expert teleoperator monitors the task’s progress, makes real-time movement commands via a SpaceMouse, and adjusts the admittance gains using keyboard input to switch between the following predefined gain modes: low-gain mode, high-gain mode, insertion mode, and contact mode.
-The low/high gain mode has low/high gains in all directions, the insertion mode has high gains in the $z$ direction of the end-effector frame and low gains elsewhere.
-Finally, the contact mode has low gains in the $z$ direction and high gains elsewhere.
-We collected $86$ demonstrations to train a near-perfect policy.
-Since we know the fixed pose of the platform a priori, e.g., a ground-truth reference frame, the GCEV vector can be calculated for the training process.
-Nevertheless, the reference frame needs to be estimated via Diff-EDF (as $g_{{}_{EDF}}$) during the inference stage, which may have non-negligible errors.
-To handle this issue, we have added noise to the reference frame $g_{ref}$ to calculate $e_{{}_{G}}$ during dataset preprocessing.
-This provides the model with an inductive bias to primarily rely on $e_{{}_{G}}$ values for rough alignment and rely on vision feedback for fine-grained motion.
-The rest of the training follows the standard imitation learning pipeline.
-IV-B2 Geometric Admittance Control (GAC)
+The G-CompACT $\pi_{\theta}$ represented in the spatial frame satisfies the following equivariance property:
+$\begin{split}&(g_{l}g_{d},K_{p},K_{R})(k)=\pi_{\theta}(g_{l}\circ o(k)).\end{split}$ (7)
+The proof is presented in Appendix AI.
+IV-B Geometric Admittance Control (GAC)
 We implement the geometric impedance control (GIC) proposed in [20, 21] in the geometric admittance control (GAC) setup [19].
@@
 where $K_{p},K_{R}\in\mathbb{R}^{3\times 3}$ symmetric being positive stiffness matrices for the translational and rotational dynamics, respectively.
-Following the standard admittance control formulation, the desired end-effector pose command is calculated using (9), which is passed to the lowest level tracking controller.
+The desired end-effector pose command is calculated using (8), which is then passed to the robot as the pose command signal.
 For details on GIC/GAC, we refer readers to [20, 19].
-The admittance control loop is implemented at a $200\mathrm{Hz}$ rate using ROS2.
-IV-C EquiContact
-The proposed EquiContact method comprises the high-level Diff-EDF, which serves as a high-level vision planner that provides the reference frame to be fed to GCEV, and the low-level G-CompACT, which handles fine-grained movement and contact interaction during insertion using real-time vision and force feedback.
-The overall pipeline of the EquiContact is presented in Fig.
-1, and also summarized in Algorithm.
-1.
-In Proposition 2, we demonstrate that the proposed EquiContact possesses the equivariance property.
+IV-C Diffusion-Equivariant Descriptor Field (Diff-EDF)
+Diffusion-Equivariant Descriptor Field (Diff-EDF) [18] is an $SE(3)$-equivariant reference frame estimator for pick-and-place tasks.
+In EquiContact, Diff-EDF serves as a high-level vision module that provides a coarse target reference frame for the downstream localized policy.
+Given a scene point cloud $\mathcal{O}^{scene}$ and a gripper point cloud $\mathcal{O}^{grasp}$ expressed in the end-effector frame, Diff-EDF outputs an estimated target pose $g_{{}_{EDF}}\in SE(3)$:
+$g_{{}_{EDF}}=f_{\varphi}(\mathcal{O}^{scene},\mathcal{O}^{grasp}).$ (10)
+Diff-EDF is designed to be left-equivariant with respect to $SE(3)$ transformations of the target object [18].
+Let $\mathcal{O}^{ref}\subset\mathcal{O}^{scene}$ denote the subset of points corresponding to the object of interest, e.g., hole assembly.
+Then, for any $g_{l}\in SE(3)$,
+$f_{\varphi}(g_{l}\circ\mathcal{O}^{ref},\mathcal{O}^{grasp})=g_{l}\cdot f_{\varphi}(\mathcal{O}^{ref},\mathcal{O}^{grasp}).$ (11)
+EquiContact relies only on the equivariance property of the reference frame estimator; the specific architecture of Diff-EDF is otherwise not essential and may be replaced by any $SE(3)$-equivariant reference frame estimator.
+Importantly, the localized policy G-CompACT is left-invariant by construction and does not require an $SE(3)$-equivariant estimator.
+In the absence of an equivariant reference frame estimator, the overall pipeline no longer guarantees end-to-end $SE(3)$ equivariance; still, the local equivariance of G-CompACT is preserved.
+IV-D EquiContact
+The proposed EquiContact method comprises the high-level Diff-EDF and the low-level G-CompACT.
+In Proposition 2, we demonstrate that if an $SE(3)$ equivariant reference frame estimator, such as Diff-EDF, is used, then the resulting EquiContact possesses the equivariance property.
 Let EquiContact be written as $h_{\Theta}$ so that $h_{\Theta}(g,g_{ref},F_{e})\mapsto f_{{}_{G}}$, i.e., $h_{\Theta}:SE(3)\times SE(3)\times\mathbb{R}^{6}\to\mathbb{R}^{6}$.
 Proposition 2.
-Suppose that the Assumption 1 and 2 hold. The EquiContact Policy $h_{\theta}$ is equivariant if it is described relative to the spatial frame.
-The proof is shown in the Appendix. -B.
+Suppose that the Assumption 1 holds. The EquiContact policy $h_{\Theta}$ is equivariant if it is described relative to the spatial frame.
+The proof is shown in the Appendix AI.
+IV-E Extensions to Pick Tasks
+So far, we have described our method in terms of the insertion (placement) task.
+The proposed method can be extended to pick tasks in the same manner.
+The Diff-EDF can be utilized to obtain the pick reference frame, which is used for $e_{{}_{G}}$ for the picking G-CompACT.
+The picking G-CompACT is trained in such a way that the manipulator grasps a peg in a fixed, aligned pose, which helps EquiContact bypass the right-equivariance issue.
+For G-CompACT, the FT sensor values are not utilized as one of its observations, and it does not output the admittance gains; instead, it uses fixed gains.
+V Experiments and Discussions
+We have conducted sets of experiments to validate the proposed $SE(3)$ vision-to-force equivariance property of the EquiContact. In particular, we aim to answer the following research questions:
+  • RQ1
+    What are the key principles for spatially generalizable contact-rich manipulation tasks?
+  • RQ2
+    Can the EquiContact framework be extended to general contact-rich tasks other than the PiH task?
+  • RQ3
+    Do our design choices of EquiContact really lead to spatial generalization?
+First, to answer RQ1, we compare the proposed EquiContact against three baselines in the PiH task: ACT with world frame observations and actions, executed with and without GAC, and CompACT [8].
+To answer RQ2, we have trained and tested EquiContact with known reference frames on the screwing and surface wiping tasks – See Fig.
+6.
+Finally, to answer RQ3, we have tested EquiContact with known reference frames on the extreme transformation scenarios for all tasks as shown in Fig.
+3.
+Figure 6: (Left) Screwing task is first to align the peg to the hole and screw-insert the peg.
+(Right) Surface wiping (erasing) task is to erase the black marker lines with the eraser.
+The same platform structure of the PiH is used.
+Before diving into the experimental results, we introduce the implementation details for training and inference of EquiContact.
+V-A Implementation details
+Training: First, we note that the G-CompACT and Diff-EDF are separately trained but executed in a single pipeline.
+To train a G-CompACT, we collect expert demonstrations via teleoperation at a fixed platform pose.
+We have collected a dataset not only with a pure white background but also with arbitrary visual distractors, so that the policy can learn to reject background perturbations.
+We have provided 1313-$18$ prompts for each task.
+The details regarding the gain modes, language prompts, and scene randomization are provided in Appendix AII.
+Since we know the platform’s fixed pose a priori during training, e.g., a ground-truth reference frame, the GCEV vector can be computed.
+Nevertheless, the reference frame needs to be estimated via Diff-EDF (as $g_{{}_{EDF}}$) during the inference stage, which may have non-negligible errors.
+To handle this issue, we have added noise to the reference frame $g_{ref}$ to calculate $e_{{}_{G}}$ during dataset preprocessing.
+This provides the model with an inductive bias to primarily rely on $e_{{}_{G}}$ values for rough alignment and rely on vision feedback for fine-grained motion.
+The rest of the training follows the standard imitation learning pipeline.
+To train Diff-EDF, the scene and grasp point clouds are collected together with the target reference frames, which represent the desired poses of the end-effector for pick-and-place operations.
+$20$ demonstrations were collected for the Diff-EDF: $10$ samples of the flat platform and $10$ samples of the tilted platform, both translationally and rotationally randomized.
+The training process of Diff-EDF follows the procedure in [18].
+Inference: We have implemented the EquiContact pipeline using the ROS2 framework.
+First, the scene point clouds are obtained and processed by Diff-EDF, producing reference frames for pick-and-place.
+Using these reference frames, the robot first moves near them, and the G-CompACT is activated near the target objects.
+During inference, we first obtain the task tokens from the previously used 131813-18 task prompts and feed the mean value of these tokens to the policy.
+The overall pipeline of the EquiContact is presented in Fig.
+LABEL:fig:equi_contact, and also summarized in Algorithm.
+1 in the Appendix.
+Please refer to the Appendix AII for in-depth implementation details.
+V-B Peg-in-Hole Benchmark Results
+Table. I summarizes the observation/action representations used in each method and reports the benchmark results across all setups.
+TABLE I: Success rates of the insertion policies in real-world experiments for the proposed and benchmark approaches.
+“In-Dist.” denotes in-distribution data and “OOD” denotes out-of-distribution data.
+For the In-Dist.
+(in distribution) scenario, the initial pose of the end-effector is randomized around the flat platform.
+Methods Observation Action Test Scenario Success Rate
+ACT w/o GAC [World Pose] [World Pose] Flat Platform (In-Dist.) 1 / 5
+ACT w/ GAC [World Pose] [World Pose] Flat Platform (In-Dist.) 18 / 20
+CompACT [World Pose, FT] [World Pose, Gains] Flat Platform (In-Dist.) 19 / 20
+Flat Platform (OOD) 0 / 10
+EquiContact [GCEV, FT] [Relative Pose, Gains] Flat Platform (OOD) 20 / 20
+(Place, Ours) Tilted Platform ($30^{\circ}$, OOD) 19 / 20
+V-B1 Demonstration of Compliance
+As the importance of left-invariant compliant control action has already been verified in [19], we focus on verifying the necessity of compliant control action.
+We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC).
+Results for this comparison are shown in the 1st and 2nd rows of Table I.
+Without GAC, the ACT model shows significantly lower success rates.
+The failure mode of the ACT w/o GAC involves collision: as the robot approaches the platform, excessive contact forces trigger safety shutdowns, preventing the task from completing.
+Due to the safety issue, we limited the number of trials without GAC to $5$.
+This result demonstrates that compliant control is nearly a deciding factor between success and failure in contact-rich tasks.
+The advantage of variable-compliant gain is that one can achieve the desired force interaction behavior through admittance gains.
+To show this, we compare ACT with GAC and CompACT under in-distribution (In-Dist.) flat platform settings.
+Although both methods achieve near-perfect success rates, their force profiles during insertion differ substantially.
+As shown in Fig.
+7, CompACT, which outputs task-adaptive admittance gains based on force-torque feedback, consistently produces lower interaction forces, especially in the $z$-direction.
+Note that during the data collection, we modulate the gains to reduce the force interaction in the $z$-direction but do not consider the magnitudes of torques.
+As a result, CompACT showed higher interaction torque throughout the task.
+The effectiveness of the CompACT compared to the baseline ACT was already presented in [8].
+Figure 7: Force profiles of CompACT and ACT with GAC (fixed gains) during insertion tasks are presented.
+The CompACT with force-torque sensor inputs and output gains shows lower interaction force in all directions.
+V-B2 Demonstration of Equivariance
+Although the CompACT succeeds in insertion tasks in trained scenarios without excessive force exertion, it fails to generalize to spatially unseen configurations.
+This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance.
+The result of applying CompACT to the translationally unseen cases is shown in the Table I 33rd row - Flat Platform (OOD).
+Note that only a flat platform is used, meaning it is randomized only translationally.
+We tested for $10$ cases and did not conduct more tests because it resulted in a $0\%$ success rate.
+In contrast, the proposed method (EquiContact) achieves perfect success rates on the translationally unseen flat platform, as can be seen in the $4$th row of Table I.
+As shown in Section IV, EquiContact has $SE(3)$ vision-to-force equivariance, achieving a near-perfect success rate, even on the tilted platform, which undergoes a full $SE(3)$ transformation.
+We attribute the single failure case to a large error from Diff-EDFs that exceeded the noise level applied during training.
+TABLE II: Success Rates of the proposed EquiContact for a full pipeline of pick-and-place.
+Test Scenario Success Rate Failure Cases
+Flat Platform (OOD) 20 / 20 N/A
+Tilted Platform ($30^{\circ}$, OOD) 19 / 20 1 Place
+The result of EquiContact for the full pick-and-place task is summarized in Table II.
+The EquiContact also demonstrates a near-perfect success rate in the full pick-and-place pipeline for peg-in-hole tasks.
+V-C Validating Feasibility to Other Contact-rich Tasks
+To further validate the EquiContact framework, we test it on two additional contact-rich tasks: screwing and surface wiping (see Fig.
+6).
+In the screwing task, the robot aligns and screws a peg; in the wiping task, it erases a line from a whiteboard.
+In this experiment, we assume that the reference frames are known.
+The reference frame for screwing is the end-effector pose at full insertion; for wiping, it is the center of the board.
+As in the PiH setup, demonstrations are collected on a fixed platform and evaluated on out-of-distribution configurations, including tilted platforms.
+Results in Table III show consistent success rates across all conditions, confirming EquiContact’s feasibility for various contact-rich tasks.
+TABLE III: Success Rates of the G-CompACT for screwing and surface wiping tasks. The evaluation is conducted with the ground-truth reference frames.
+Test Scenario Screwing Wiping
+Flat Platform (OOD) 10 / 10 10 / 10
+Tilted Platform ($30^{\circ}$, OOD) 9 / 10 10 / 10
+V-D Results on Extreme Transformation Cases
+TABLE IV: Success Rates of the G-CompACT for PiH, screwing, and surface wiping tasks to extreme task transformations (See Fig. 3). The evaluation is conducted with the ground-truth reference frames.
+Testing Scenarios PiH Screwing Wiping
+$45^{\circ}$ in $y$ 10 / 10 4 / 10 10 / 10
+$90^{\circ}$ in $x$ 9 / 10 0 / 10 0 / 5
+$135^{\circ}$ in $x$ 10 / 10 2 / 10 0 / 5
+We test the EquiContact method on the extreme transformed configuration for each task.
+The results are summarized in Table IV.
+As described in Fig.
+5, extreme task transformation cases have different camera inputs, e.g., unseen background objects and light & shadow variations.
+Therefore, the requirement for invariant visual features becomes more stringent, and a robust vision encoder is needed.
+The G-CompACT for the PiH task showed almost perfect success rates on all extreme task transformations, validating that a well-trained G-CompACT policy can handle arbitrary task transformations.
+For the surface wiping tasks, the trained G-CompACT policy can handle the smallest angle perturbation $45^{\circ}$, but failed completely on the other cases.
+The failure mode is that the robot is trying to track the slots of aluminum extrusion or black cables, not the black lines on the whiteboard.
+This may be because the prompts we provided for surface wiping include the phrase “the black line.” In addition, the lack of a prominent blue square mark on the target object might be the issue.
+In order to overcome this, one might need to provide more diverse visual distractors that are similar to the black lines, so that the transformer can learn meaningful cross-attention between the vision and GCEV signals.
+We have not tried more than $5$ trials, as it consistently fails.
+For the screwing task, we have relaxed the success condition to insert and rotate by at least $20\%$ due to vibration.
+Unlike the flat platform and $30^{\circ}$ platform cases, where the platforms are tightly assembled on the optical table, the platforms for extreme transformation cases are attached to the aluminum extrusion cages, as in Fig.
+3.
+However, as the aluminum extrusion cages are cantilever beams fixed to the optical tables, exerting forces in $x$ and $y$ directions on the end-effector frame leads to high vibration, resulting in complete failures.
+Despite the relaxed success criterion, the success rates of screwing to these scenarios are significantly lower than those of mild transformations.
+This is because the screwing task is much more complicated than the PiH task, since it requires perfect alignment to be inserted and progressed.
+Therefore, we might need a more diverse and carefully curated dataset to finish the task on these cases successfully.
+V-E Limitations and Future Work
+Symmetry Braking: The most prominent failure case of EquiContact is the symmetry braking, specifically, manipulator singularities.
+When the robot is near singular, pose tracking accuracy degrades because controllers sacrifice tracking to avoid singularity, leading to poorly executed policy commands, resulting in a distributional shift issue.
+Therefore, the testing scenarios for extreme transformations are carefully selected so that the singularities are not encountered during execution.
+Lack of Dataset: Especially for the surface wiping and screwing tasks, the overall quality of the policy could be increased with a larger dataset of better quality.
+For the surface wiping, the variations of the backgrounds need to be increased, and for the screwing, the demonstration examples with less force interaction are required.
+In the latter case, a teleoperation device that provides force feedback, i.e., a bilateral teleoperation [10] device, would be beneficial.
+Generalization to Other Visuomotor Policies: We have utilized an ACT-based visuomotor policy in our current work, but our EquiContact framework can be generalized to diffusion policy (DP) or flow-matching style.
+Vision Encoders: We employed a language-guided visual feature to realize a left-invariance.
+In fact, the left-invariant vision encoder is closely related to the object-centric representation, such as slot attentions [11].
+Moreover, although we have used CLIP-RN50 as our vision backbone, newer versions of vision-language models (VLMs) are available, such as SigLIP [35].
+Notably, recent works [12, 31] explored inducing 3D equivariance from 2D images.
+We will investigate these works of vision encoders to improve left-invariance for future work.
+VI Conclusion
+In this work, we introduced EquiContact, a vision-to-force equivariant policy for spatially generalizable contact-rich tasks.
+By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations.
+We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations, validated its effectiveness through real-world experiments on PiH benchmarks, and its feasibility towards screwing and surface wiping tasks.
+Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates.
+Through extensive benchmark studies, we highlighted the effectiveness of the three principles – compliance, localized policy, and induced equivariance – for achieving spatial generalizability in contact-rich manipulation.
+We conclude that these principles offer a simple yet powerful design guideline for developing spatially generalizable and interpretable robotic policies complementing recent trends in end-to-end visuomotor learning and enabling a structured divide-and-conquer approach.
+Appendix
+AI Proof of Propositions
+In this section, we present the detailed proofs omitted from the main manuscript. We begin by introducing some preliminaries.
+AI-A Preliminaries
+We are interested in the matrix Lie group representation $g$ of the manipulator’s end-effector pose, where $g\in SE(3)$, given by
+$g=\begin{bmatrix}R&p\\ 0&1\end{bmatrix}\in SE(3),$ (12)
+where $SE(3)$ is a Special Euclidean group, $R\in SO(3)$ with $SO(3)$ being a Special Orthogonal group, and $p\in\mathbb{R}^{3}$. We first define invariance and equivariance.
+Definition 1 ($SE(3)$ left invariance and equivariance).
+Let $f$ be a function $f:\mathcal{X}\to\mathcal{Y}$, so that $y=f(x)$, where $x\in\mathcal{X}$ is a domain and $y\in\mathcal{Y}$ is a co-domain.
+Then, a function $f$ is left-invariant to $SE(3)$ (left) group action $g_{l}\in SE(3)$ if the following equation holds:
+$f(g_{l}\circ x)=f(x),$ (13)
+where $\circ$ is a group action on the domain or co-domain .
+Similarly, a function $f$ is left-equivariant to $SE(3)$ group action $g_{l}$ if the following holds:
+$f(g_{l}\circ x)=g_{l}\circ f(x).$ (14)
+In fact, the group action is realized by the appropriate group representation on the acting set, i.e., the domain or co-domain, which is often denoted by $\rho(g_{l})$ in group equivariant deep learning literature [22].
+Important examples widely used throughout the paper include cases where the domain or co-domain is $SE(3)$ itself or the wrench [^1] Although the original wrench should be represented in $se^{*}(3)$, a dual-space of Lie-algebra, we use a vector representation of $se^{*}(3)$ to reduce mathematical details.
+$\mathbb{R}^{6}$.
+If the set that the group acts on is the $SE(3)$ group itself, then,
+$g_{l}\circ g=g_{l}\cdot g,\quad\forall g,g_{l}\in SE(3),$ (15)
+where $\cdot$ is a standard matrix multiplication. If the set that the group actions on is the wrench $\mathbb{R}^{6}$, then,
+$g_{l}\circ h=\text{Ad}_{g_{l}^{\text{-}1}}^{T}h,\quad\forall g_{l}\in SE(3),\;\forall h\in\mathbb{R}^{6},$ (16)
+where $\text{Ad}:SE(3)\times\mathbb{R}^{6}\to\mathbb{R}^{6}$, defined as
+$\text{Ad}_{g_{l}}=\begin{bmatrix}R_{l}&\hat{p}_{l}R_{l}\\ 0&R_{l}\end{bmatrix}\in\mathbb{R}^{6\times 6},$ (17)
+with $g_{l}=(p_{l},R_{l})$. For the details of the group action and representation, we refer to [22].
+Lemma 1 (Left invariance of GCEV and elastic wrench [19]).
+The GCEV $e_{{}_{G}}$ (2) and elastic wrench $f_{{}_{G}}$ (9) is left-invariant, i.e., $\forall g_{1},g_{2},g_{l}\in SE(3)$,
+$\begin{split}e_{{}_{G}}(g_{l}\circ(g_{1},g_{2}))&=e_{{}_{G}}(g_{l}g_{1},g_{l}g_{2})=e_{{}_{G}}(g_{1},g_{2}),\\ f_{{}_{G}}(g_{l}\circ(g_{1},g_{2},K_{p},K_{R}))&=f_{{}_{G}}(g_{l}g_{1},g_{l}g_{2},K_{p},K_{R})\\ &=f_{{}_{G}}(g_{1},g_{2},K_{p},K_{R}).\end{split}$ (18)
+Proof.
+Although the full proof is presented in [19], we include it for completeness.
+Let $g_{l}=(p_{l},R_{l})$, and $g_{i}=(p_{i},R_{i})$ with $i=\{1,2\}$.
+Then, the left-transformed homogeneous matrix $g_{l}g_{i}$ is calculated in the following way:
+$g_{l}g_{i}=\begin{bmatrix}R_{l}&p_{l}\\ 0&1\end{bmatrix}\begin{bmatrix}R_{i}&p_{i}\\ 0&1\end{bmatrix}=\begin{bmatrix}R_{l}R_{i}&R_{l}p_{i}+p_{l}\\ 0&1\end{bmatrix}$,
+i.e., $g_{l}g_{i}=(R_{l}p_{i}+p_{l},R_{l}R_{i})$ The left-transformed GCEV is then
+$e_{{}_{G}}(g_{l}g_{1},g_{l}g_{2})$ $=\begin{bmatrix}R_{1}^{T}R_{l}^{T}\left(R_{l}p_{1}+p_{l}-R_{l}p_{2}-p_{l}\right)\\ \left((R_{l}R_{2})^{T}R_{l}R_{1}-(R_{l}R_{1})^{T}R_{l}R_{2}\right)^{\vee}\end{bmatrix}$
+$=\begin{bmatrix}R_{1}^{T}(p_{1}-p_{2})\\ (R_{2}^{T}R_{1}-R_{1}^{T}R_{2})^{\vee}\end{bmatrix}=e_{{}_{G}}(g_{1},g_{2}),$ (19)
+where the definition of the rotation matrix is used, i.e., $R^{T}R=RR^{T}=I$, $\forall R\in SO(3)$. Similarly, the left-transformed elastic wrench reads
+$f_{{}_{G}}(g_{l}g_{1},g_{l}g_{2},K_{p},K_{R})$
+$=\begin{bmatrix}(R_{l}R_{1})^{T}R_{l}R_{2}K_{p}(R_{l}R_{2})^{T}(R_{l}p_{1}+p_{l}-R_{l}p_{2}-p_{l})\\ (K_{R}(R_{l}R_{2})^{T}R_{l}R_{1}-(R_{l}R_{1})^{T}R_{l}R_{2}K_{R})^{\vee}\end{bmatrix}$
+$=\begin{bmatrix}R_{1}^{T}R_{2}K_{p}R_{2}(p_{1}-p_{2})\\ (K_{R}R_{2}^{T}R_{1}-R_{1}^{T}R_{2}K_{R})^{\vee}\end{bmatrix}=f_{{}_{G}}(g_{1},g_{2},\!K_{p},\!K_{R}).$ (20)
+We note that the gains $K_{p}$ and $K_{R}$ are defined on the desired frame [2], i.e., on the body-frame; therefore, they are not affected by left-group actions (change of spatial coordinate system).
+AI-B Proof of Proposition 1
+The left-transformed observation signals $g_{l}\circ o(k)$ reads that:
+$g_{l}\circ o(k)=(g_{l}\circ e_{{}_{G}},g_{l}\circ F_{e},g_{l}\circ I_{w}).$ (21)
+As was shown in Lemma 1, the GCEV $e_{{}_{G}}$ is left invariant as
+$g_{l}\circ e_{{}_{G}}(g,g_{{}_{EDF}})=e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}})=e_{{}_{G}}(g,g_{{}_{EDF}}).$
+The force-torque sensor values are left-invariant because they are already defined with respect to the end-effector frame [19], and the visual representation vectors satisfy left invariance if Assumption 1 is ideally met.
+Combining all these properties, it follows that
+$a(k)$ $=\pi_{\theta}(g_{l}\circ o(k))=\mathcal{D}_{\psi}(g_{l}\circ e_{{}_{G}},g_{l}\circ F_{e},\mu_{\phi}(g_{l}\circ I_{w}))$
+$=\mathcal{D}_{\psi}(e_{{}_{G}},F_{e},z)=\pi_{\theta}(o(k)),$ (22)
+which shows the left invariance of the G-CompACT policy on the end-effector frame. $\blacksquare$
+AI-C Proof of Corollary 1
+Notice that the desired pose signal in the spatial frame $g_{d}$ is obtained via (with a slight abuse of notation)
+$g_{d}=gg_{rel}=g\cdot\pi_{\theta}(o(k))\triangleq\pi_{\theta}^{s}(o(k)),$ (23)
+where the superscript $s$ denotes that the policy is described on the spatial frame, i.e., the world frame. Then, utilizing the left-invariance property from Proposition 1, it follows that
+$\begin{split}(g_{l}g_{d},K_{p},K_{R})&=\pi^{s}_{\theta}(g_{l}\circ o(k))\\ &=g_{l}g\cdot\pi(g_{l}\circ o(k))=g_{l}g\cdot\pi(o(k)).\end{split}$ (24)
+Therefore, when the policy is left-transformed by an arbitrary element $g_{l}\in SE(3)$, the resulting trajectories in the spatial frame $g_{d}$ are also transformed to $g_{l}g_{d}$, showing the equivariance property.
+$\blacksquare$
+AI-D Proof of Proposition 2
+Let the object of interest, e.g., a peg for the picking task and a hole for the placing task, be observed by $\mathcal{O}^{ref}$ and $I_{w}$ with its pose given by $g_{ref}$, so that the left-translated $g_{l}\cdot g_{ref}$ is observed by $g_{l}\circ\mathcal{O}^{ref}$ from the point cloud, and $g_{l}\circ I_{w}$ by the left-translated end-effector attached wrist camera as described in Fig.
+5.
+First, notice that $h_{\Theta}$ can be fully written as
+$\begin{split}h_{\Theta}(g,g_{ref},F_{e})&=f_{{}_{G}}(g,g_{d},K_{p},K_{R})\\ &=f_{{}_{G}}(g,{\pi_{\theta}^{s}(e_{{}_{G}}(g,g_{{}_{EDF}}),F_{e},I_{w})})\\ &=f_{{}_{G}}(g,\pi_{\theta}^{s}(e_{{}_{G}}(g,f_{\varphi}(\mathcal{O}^{ref})),F_{e},I_{w}))\end{split}$ (25)
+Note also that $g_{{}_{EDF}}$ is fed to $e_{{}_{G}}$, not $g_{ref}$.
+Then, when both $g$ and $g_{ref}$ undergo a left transformation $g_{l}$, from Assumption 1 and Corollary 1, the following holds:
+$h_{\Theta}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})$
+$=f_{{}_{G}}(g_{l}g,\pi_{\theta}^{s}(e_{{}_{G}}(g_{l}g,f_{\varphi}(g_{l}\circ\mathcal{O}^{ref})),g_{l}\circ F_{e},g_{l}\circ I_{w}))$
+$=f_{{}_{G}}(g_{l}g,\pi_{\theta}^{s}(e_{{}_{G}}(g_{l}g,g_{l}g_{{}_{EDF}}),g_{l}\circ F_{e},g_{l}\circ I_{w}))$ (26)
+$=f_{{}_{G}}(g_{l}g,g_{l}g_{d},K_{p},K_{R})=f_{{}_{G}}(g,g_{d},K_{p},K_{R})$
+$=h_{\Theta}(g,g_{ref},F_{e}).$
+The second-last equation ($SE(3)$ left-invariance of the elastic wrench) comes from Lemma 1.
+Finally, as the $h_{\Theta}$ is left-invariant and is defined on the end-effector frame, from the result of Proposition 2 of [19], $h_{\Theta}$ is equivariant, if it is described in the spatial frame, i.e.,
+$h_{\Theta}^{s}(g_{l}g,g_{l}g_{ref},g_{l}\circ F_{e})=\text{Ad}_{g_{l}^{\text{-}1}}^{T}h_{\Theta}(g,g_{ref},F_{e}),$ (27)
+where Ad is a (large) adjoint operator. From Definition 1, $h_{\Theta}^{s}$ is an equivariant function [19]. $\blacksquare$
+AII Implementation Details
+AII-A G-CompACT Training
+AII-A1 Details on models
+The objective of this chapter is to highlight the details of the selected models, especially the vision encoders.
+Our G-CompACT model has a CLIP-RN50 vision backbone that is modulated by the FiLM layer from the CLIP text encoder.
+A few notable hyperparameters are summarized in the Table A1.
+TABLE A1: Hyperparameters of G-CompACT. The other hyperparameters are adapted from the ACT [37].
+Names Values
+Image Size $[224,224]$
+Learning Rate (policy) $\eta_{policy}$ 1e051e-05
+Learning Rate (Vision Encoder) $\eta_{vision}$ 1e051e-05
+Epochs $15,000$
+Batch Size $32$
+Batch Size $32$
+As denoted in Fig.
+LABEL:fig:equi_contact, we use $30\mathrm{Hz}$ of inference frequency, with the chunking size of $60$.
+In addition, we used rotation vector (rotvec) representation for relative pose actions, and 6D rotation (rot6d) representation for world-pose observation and actions for benchmark models.
+This is because the default end-effector configuration in the world (spatial) frame tends to have a $180^{\circ}$ rotation angle, leading to a sign flip when using the rotation vector representation.
+AII-A2 Dataset details
+To collect the demonstration dataset, the expert teleoperator monitors the task’s progress, makes real-time movement commands via a SpaceMouse, and adjusts the admittance gains using keyboard input to switch between predefined gain modes: low-gain mode, high-gain mode, insertion mode, and contact mode.
+Admittance Gains
+The gain modes utilized during the training are low-gain mode, high-gain mode, insertion mode, and contact mode.
+The low/high gain mode has low/high gains in all directions, the insertion mode has high gains in the $z$ direction of the end-effector frame and low gains elsewhere.
+Finally, the contact mode has low gains in the $z$ direction and high gains elsewhere.
+In our EquiContact implementation, we use $M=0.5I_{6\times 6}$.
+We used only the diagonal terms of the stiffness matrices $K_{p}$ and $K_{R}$ for learning and GAC implementation [^2] This is also a great benefit of using geometric impedance/admittance control [20]..
+Therefore, the stiffness gains $(K_{p},K_{R})$ can be represented with $6$ dimension.
+The details are as follows:
+  • Low-gain: $(K_{p},K_{R})=(300,300,300,300,300,300)$
+  • High-gain: $(K_{p},K_{R})=(1000,1000,1000,1000,1000,1000)$
+  • Contact mode: $(K_{p},K_{R})=(1500,1500,300,1500,1500,1500)$
+  • Insertion mode: $(K_{p},K_{R})=(300,300,1500,300,300,300)$
+We note that the detailed gains implementation may vary significantly depending on the specific implementation, such as sampling frequency and even robot firmware version.
+The sets of working gains are found through trial and error.
+Data Collection Methods
+For the PiH task, the end-effector is first aligned with high-gain mode, quickly converted to the contact mode to make a surface contact and search for a hole, and the insertion mode is activated when the peg is slightly inserted into the hole.
+For the surface wiping task, the high-gain mode is used to align with the whiteboard, and the surface contact mode is used during the surface wiping.
+For the screwing task, the high-gain mode is first used to align with the screw hole, followed by contact mode for fine searching.
+Then, the insertion mode is activated to ensure screw-locking, and the low-gain mode is used for screw rotation.
+Scene Randomization and details
+The initial pose of the end-effector is randomized for both with and without background variations.
+We add arbitrary lab objects with random poses for the background variations.
+The examples of the scene randomization for PiH are presented in Fig.
+A1.
+Figure A1: Examples of scene randomization during data collection are shown.
+Notice the initial pose randomization of the end-effector for all cases.
+(a) Example without background variations.
+(b)-(d) Examples with background variations composed of arbitrary lab objects with random poses.
+AII-A3 Text Prompts
+Here, we additionally provide the full text prompts used during the training.
+As mentioned in Sec.
+IV, the average text tokens are fed during the inference phase.
+We provide the full text prompts for the PiH-placing task, but show few samples for the other tasks to reduce verbosity.
+PiH - Placing
+Below is the full sets of text prompts utilized for the PiH placing task.
+  • “A yellow peg approaching a light-gray assembly target marked with blue.”
+  • “A cylindrical yellow peg inserted into a round light-gray hole on a square board with blue tape edges.”
+  • “A yellow dowel being aligned with a light-gray target that has blue markings on a flat surface.”
+  • “A yellow peg going into a light-gray round hole with blue tape border.”
+  • “A yellow cylindrical peg and a round light-gray hole with blue tape around the square board.”
+  • “A yellow plastic dowel aligning with a round light-gray socket, blue tape square.”
+  • “A thick yellow pin approaching a light-gray round socket with blue tape.”
+  • “A yellow peg being inserted into a light-gray circular hole on a board with blue tape.”
+  • “A thick yellow stick above a light-gray round socket with blue tape.”
+  • “A thick light-yellow stick aligning to a light-gray assembly target with blue marks.”
+  • “A yellow plastic dowel being inserted into a round socket with blue tape on a flat surface.”
+  • “Peg-in-hole task: yellow plastic peg and light-gray round hole with blue tape around the board.”
+  • “A cylindrical yellow peg entering a light-gray circular recess with a blue tape border.”
+  • “A yellow peg being aligned with a light-gray target that has blue markings on a flat surface.”
+  • “A yellow peg approaching a light-gray assembly target marked with blue.”
+  • “A yellow cylindrical peg inserted into a round light-gray hole on a square board with blue tape edges.”
+  • “A yellow peg reoriented to align with a light-gray circular hole bordered with blue tape.”
+  • “A yellow cylindrical peg re-aligning to fit into a round light-gray hole with blue tape edges.”
+PiH - Picking
+Among the $13$ prompts, we present $3$ prompts for example in this paper.
+  • “A black robotic gripper is about to pick up a yellow peg.”
+  • “A robotic pick-up: a black gripper and a yellow peg object.”
+  • “A black robotic gripper reaching toward a yellow cylindrical dowel”
+Screwing Task
+Below is the task prompts example of screwing task among $16$ prompts.
+  • “A yellow cylindrical peg rotating inside a round light-gray hole with blue tape border.”
+  • “yellow plastic dowel aligning with a round light-gray socket, blue tape square.”
+  • “A yellow peg being aligned and screwed into a light-gray target with blue markings.”
+Surface Wiping Task
+Below is the task prompts example of surface wiping task among $16$ prompts.
+  • “A black metallic robotic gripper wiping black markings with a yellow eraser.”
+  • “robotic gripper grasping a yellow eraser moving on top of the black markings.”
+  • “a black robot gripper holding a yellow eraser moving over black lines.”
+  • “black markings being erased by a yellow eraser held by a robot gripper.”
+AII-B Diff-EDF Implementation Details
+Instead of following the original Diff-EDF pipeline, which utilizes pick-and-place models, we used two pick models.
+This decision mainly stems from the task setup, where the peg is upright, and the peg is grasped by the gripper in an aligned, upright pose.
+The core difference of the pick and place model is that the place model needs to get the grasp point cloud after each grasp to handle the right equivariance of the model.
+However, from task setup, we bypass this right equivariance issue, removing the necessity of the place model.
+We also used the post-processing heuristics to filter the output pose of the Diff-EDF.
+The Diff-EDF outputs $20$ candidate target poses for picking and $20$ candidates for the place, which are ranked by the energy level.
+Although in theory, the lower energy poses should result in a better pose, we found out that this does not hold in practice.
+Instead, we figured out that the Diff-EDF have low variations on the position of the “tip”.
+The mean value of the tip position candidates are first calculated and is used to recalculate the pose of the end-effector from the known orientation (upright peg assumption).
+On the other hand, for the placing, the desired orientation is calculated by taking a mean value of the orientation.
+The position of the end-effector is similarly calculated from the position of the tip.
+AII-C GAC Implementation
+We implement the geometric admittance controller (GAC) using the pose tracking controller.
+Given the desired dynamics (8), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
+$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
+$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (28)
+where $T_{s}$ is a sampling time ($5\mathrm{ms}$ for GAC) and $\hat{(\cdot)}$ denotes a hat-map.
+AII-D Full Pipeline Implementation
 Algorithm 1 Inference Procedure of EquiContact
@@
 2:Get scene and grasp point cloud $\mathcal{O}^{scene}$, $\mathcal{O}^{grasp}$
-3:Run Diff-EDF for reference frame $g_{{}_{EDF}}\!=\!f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$
+3:Run Diff-EDF for reference frame $g_{{}_{EDF}}\!=\!f_{\varphi}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$
 4:Move the end-effector near the reference frame and
@@
 9:  Run G-CompACT:
-10:      $(g_{rel},K_{p},K_{R})(k)=\pi_{\theta_{2}}(e_{{}_{G}},F_{e},I_{w})(k)$ (3)
+10:      $(g_{rel},K_{p},K_{R})(k)=\pi_{\theta}(e_{{}_{G}},F_{e},I_{w})(k)$ (1)
 11:  Calculate desired pose $g_{d}(k)=g(k)g_{rel}(k)$
@@
 14:end for
-IV-D Extensions to Pick Tasks
-So far, we have described our method in terms of the insertion (placement) task.
-The proposed method can be extended to pick tasks in the same manner.
-The Diff-EDF can be utilized to obtain the pick reference frame, which is used for $e_{{}_{G}}$ for the picking G-CompACT.
-The picking G-CompACT is trained in such a way that the manipulator grasps a peg in a fixed, aligned pose, which helps EquiContact bypass the right-equivariance issue.
-For G-CompACT, the FT sensor values are not utilized as one of its observations, and it does not output the admittance gains; instead, it uses fixed gains.
-V Experimental Results and Discussions
-V-A Benchmark Results
-To evaluate the effectiveness of EquiContact and its core design principles, we compare it against three baselines: ACT with world-frame observations and actions, executed with and without GAC, and CompACT [17].
-Note that GAC reduces to standard admittance control when the platform is fixed, i.e., not tilted.
-Table.
-I summarizes the observation/action representations used in each method and reports the benchmark results across all setups.
-TABLE I: Success rates of the insertion policies in real-world experiments for the proposed and benchmark approaches.
-“In-Dist.” denotes in-distribution data and “OOD” denotes out-of-distribution data.
-For the In-Dist.
-(in distribution) scenario, the initial pose of the end-effector is randomized around the flat platform.
-Methods Observation Action Cameras Test Scenario Success Rate
-ACT w/o GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 2 / 10
-ACT w/ GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
-CompACT [World Pose, FT] [World Pose, Gains] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
-Flat Platform (OOD) 0 / 10
-EquiContact [GCEV, FT] [Relative Pose, Gains] 2 Wrists Cameras Flat Platform (OOD) 20 / 20
-(Place, Ours) Tilted Platform ($30^{\circ}$, OOD) 19 / 20
-V-A1 Demonstration of Compliance
-As the importance of left-invariant compliant control action has already been verified in [11], we focus on verifying the necessity of compliant control action.
-We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC).
-Results for this comparison are shown in the 1st and 2nd rows of Table I.
-Without GAC, the ACT model shows significantly lower success rates.
-The failure mode of the ACT w/o GAC involves collision: as the robot approaches the platform, excessive contact forces trigger safety shutdowns, preventing the task from completing.
-Due to the safety issue, we limited the number of trials without GAC to 10.
-This result demonstrates that compliant control is nearly a deciding factor between success and failure in contact-rich tasks.
-To further analyze the effects of compliant gain control, we compare ACT with GAC and CompACT under in-distribution (In-Dist.) platform settings.
-Although both methods achieve perfect success rates, their force profiles during insertion differ substantially.
-As shown in Fig.
-4, CompACT, which outputs task-adaptive admittance gains based on force-torque feedback, consistently produces lower interaction forces in all directions.
-The effectiveness of the CompACT compared to the baseline ACT was already presented in [17].
-Figure 4: Force profiles of CompACT and ACT with GAC (fixed gains) during insertion tasks are presented.
-The CompACT with force-torque sensor inputs and output gains shows lower interaction force in all directions.
-V-A2 Demonstration of Equivariance
-Although the CompACT succeeds in insertion tasks in trained scenarios without excessive force exertion, it fails to generalize to spatially unseen configurations.
-This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance.
-The result of applying CompACT to the translationally unseen cases is shown in the Table I 33rd row - Flat Platform (OOD).
-Note that the flat platform is only utilized, meaning it is only randomized translationally.
-We tested for $10$ cases and did not conduct more tests because it resulted in a $0\%$ success rate.
-In contrast, the proposed method (EquiContact) achieves perfect success rates on the translationally unseen flat platform, as can be seen in the $4$th row of Table I.
-As shown in Section IV, EquiContact has $SE(3)$ vision-to-force equivariance, achieving a near-perfect success rate, even on the tilted platform, which undergoes a full $SE(3)$ transformation.
-We attribute the single failure case to a large error from Diff-EDFs that exceeded the noise level applied during training.
-TABLE II: Success Rates of the proposed EquiContact for a full pipeline of pick and place.
-Test Scenario Success Rate Failure Cases
-Flat Platform (OOD) 20 / 20 N/A
-Tilted Platform ($30^{\circ}$, OOD) 18 / 20 1 Pick, 1 Place
-The result of EquiContact for the full pick-and-place task is summarized in Table II.
-The EquiContact also demonstrates a near-perfect success rate in the full pick-and-place pipeline for peg-in-hole tasks.
-However, as the whole task is formulated in sequential stages, the error in the previous stage tends to propagate to the following stages, leading to slightly increased failure cases.
-V-B Limitations and Discussions
-TABLE III: Success Rates of the G-CompACT (insertion) trained with the base dataset and augmented dataset for the flat platform under the presence of visual distractor scenario and tilted platform with large angle cases.
-The evaluation is conducted with the ground-truth reference frames.
-Test Scenario Base Dataset Aug. Dataset
-W/ Visual Distractor 4 / 10 9 / 10
-Tilted Platform ($45^{\circ}$) 5 / 10 9 / 10
-V-B1 Limitation on Left-invariant Visual Features
-Although EquiContact is theoretically vision-to-force equivariant, this guarantee holds only under Assumptions 1 and 2.
-In particular, Assumption 2, which requires left-invariant visual features, is challenging to ensure in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias.
-Consequently, G-CompACT shows degraded performance in settings with visual distractors or severe tilting ($45^{\circ}$), as shown in Table III (“Base Dataset”).
-The degradation arises because larger tilt angles expose unseen background regions in the wrist camera view (see Fig.
-3), resulting in out-of-distribution observations.
-This can be mitigated by augmenting the training data with visual distractors and demonstrations collected from a tilted angle.
-To demonstrate this, we added $20$ demos with distractors on a flat platform and 20 demos from a $30^{\circ}$ tilted setup.
-As shown in the “Aug.
-Dataset” column of Table III, the success rates have recovered to original levels.
-Enforcing left-invariance on RGB images via architectural design remains an open challenge.
-Recent works have explored inducing 3D equivariance from 2D images [32, 33].
-Another promising direction involves using pretrained foundation models, such as CLIP and DINO, which may exhibit implicit equivariance due to large-scale and diverse training.
-Language-guided features (e.g., using the prompt “peg”) can help extract relevant features across views.
-Alternatively, real-time segmentation models can be used to focus only on the object of interest.
-We will investigate enforcing left-invariance of images in future work.
-V-B2 Hardware and Algorithmic Limitation
-A key challenge in implementing equivariant policies for robotic manipulation is handling singularities [10].
-When the robot is near singular, pose tracking accuracy degrades because controllers sacrifice tracking to avoid singularity, leading to poorly executed policy commands.
-This results in a distributional shift, which is a critical limitation for imitation learning.
-From our experimental demonstration perspective, a small workspace of the Diff-EDF (Fig.
-2) was a drawback due to its limited receptive field.
-A combination of the singularity, joint limits, and small workspace resulted in limited demonstration of our approach, not covering a complete $SE(3)$ transformation.
-V-C Validating Feasibility to Other Contact-rich Tasks
-To further validate the EquiContact framework, we test it on two additional contact-rich tasks: screwing and surface wiping (see Fig.
-5).
-In the screwing task, the robot aligns and screws a peg; in the wiping task, it erases a line from a whiteboard.
-For simplicity, we use a white background to eliminate visual distractions and assume that reference frames are known.
-The reference frame for screwing is the end-effector pose at full insertion; for wiping, it is the center of the board.
-As in the PiH setup, demonstrations are collected on a fixed platform and evaluated on out-of-distribution configurations, including tilted platforms.
-We collected 50 demonstrations per task.
-Results in Table IV show consistent success rates across all conditions, confirming EquiContact’s spatial generalizability.
-Figure 5: (Left) Screwing task, (Right) Surface wiping (erasing) task. The same platform structure of the PiH is used.
-TABLE IV: Success Rates of the G-CompACT for screwing and surface wiping tasks. The evaluation is conducted with the ground-truth reference frames. Both tasks are trained with $50$ demos.
-Test Scenario Screwing Surface Wiping
-Flat Platform (OOD) 10 / 10 10 / 10
-Tilted Platform ($30^{\circ}$, OOD) 9 / 10 9 / 10
-VI Conclusion
-In this work, we introduced EquiContact, a vision-to-force equivariant policy for spatially generalizable contact-rich tasks.
-By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations.
-We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations, validated its effectiveness through real-world experiments on PiH benchmarks and its feasibility towards screwing and surface wiping tasks.
-Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates.
-Through extensive benchmark studies, we highlighted the effectiveness of the three principles – compliance, localized policy, and induced equivariance – for achieving spatial generalizability in contact-rich manipulation.
-We conclude that these principles offer a simple yet powerful design guideline for developing spatially generalizable and interpretable robotic policies complementing recent trends in end-to-end visuomotor learning and enabling a structured divide-and-conquer approach.
-Supplementary Material
-Implementation Details
-In this section, we provide the details of the hardware code implementation.
-Since every hardware system has a different implementation, our goal is to provide a comprehensive overview and a guide to our code distribution.
-We will use the exact names of the nodes to enable readers to understand our codebase easily.
-The hardware experimental implementation code is provided in https://github.com/Joohwan-Seo/EquiContact-Experiment, and the simulation implementation code in https://github.com/Joohwan-Seo/EquiContact-Simulation.
-S-A Implementation Pipeline
-The pipeline visualized in Fig.
-S1 is an interconnected system of camera, robot, and controller nodes with two policy services, all implemented in ROS2 Humble.
-The two policies are implemented as services because they are activated infrequently and latencies in activating the policies initially do not affect the behavior in practice.
-The execution of the full pipeline is divided into two primary stages: querying Diffusion-EDF for initial pick and insertion end-effector poses, and then executing the appropriate ACT sub-policies for these two steps after reaching a neighborhood of the predicted poses.
-In the first stage, a request to the Diffusion-EDF Service is made through the FullPipelineClient, which prompts the service to pull the individual point clouds published by the Orbbec Node.
-The service subsequently merges the point cloud in the spatial frame, preprocesses it, and passes it to the initialized Diff-EDF model to produce both pick and insertion poses.
-The Diff-EDF model, as covered in the following section, is formulated as a “double-pick” model that assumes that a fixed grip will be executed, and a cached empty gripper point cloud is passed along with the scene point cloud to get both predictions.
-In the second stage, the FullPipelineClient requests the G-CompACT Service node to complete the task with these two poses as noisy references.
-The service initializes both pick and place policies and executes the tasks serially, raising flags to signal completion of each step.
-For raw camera images and proprioceptive information, the service subscribes to nodes for each camera (from Wrist Camera Nodes) and the Robot Node respectively; note that the controller and proprioceptive helper are encapsulated in one Robot Node that manages both publishing proprioceptive information and executing policy predictions.
-During the pick portion, the pick ACT model only publishes end-effector poses, and fixed high admittance gains are utilized because the demonstrations encourage the model to make limited contact with the peg.
-The pick model has a fixed time to complete the task before a grasp is attempted, and the robot moves to the next step.
-Similarly, the place ACT policy is executed given a fixed-time limit, but for the place model, the force-torque data is also passed as an input, and admittance gains are outputted along with the end-effector poses.
-A fixed-time limit for both phases is used because it is easily implemented and works more consistently than other, more complicated completion heuristics.
-Figure S1: Whole pipeline implemented (via client named as FullPipelineClient) with ROS2 is presented.
-S-B Diffusion-EDF
-S-B1 Implementation Heuristics
-Instead of following the original Diff-EDF pipeline, which utilizes pick-and-place models, we used two pick models.
-This decision mainly stems from the task setup, where the peg is upright, and the peg is grasped by the gripper in an aligned, upright pose.
-The core difference of the pick and place model is that the place model needs to get the grasp point cloud after each grasp to handle the right equivariance of the model.
-However, from task setup, we bypass this right equivariance issue, removing the necessity of the place model.
-We also used the post-processing heuristics to filter the output pose of the Diff-EDF.
-The Diff-EDF outputs $20$ candidate target poses for picking and $10$ candidates for the place, which are ranked by the energy level.
-Although in theory, the lower energy poses should result in a better pose, we found out that this does not hold in practice.
-Instead, we figured out that the Diff-EDF have low variations on the position of the “tip”.
-To elaborate, suppose a picking policy returns the desired poses $g_{pick}^{i}=(p_{pick}^{i},R_{pick}^{i})$, which are the end-effector poses, not the tip, or Tool-Center-Point (TCP) poses.
-Then, the position of the tip estimated by Diff-EDF can be calculated via
-$p_{tip,pick}^{i}=p_{pick}^{i}+R_{pick}^{i}p_{et,pick}$
-where $p_{et,pick}\in\mathbb{R}^{3}$ denotes the coordinate of the gripper’s tip with respect to the end-effector body frame.
-Using this property and the task setup that the peg is always upright, which gives $R_{peg}$, we recalculate the Diff-EDF of the picking as
-$p_{pick}=\bar{p}_{tip,pick}-R_{peg}^{T}p_{et,pick}$
-where $\bar{p}_{tip,pick}$ is the mean of $p_{tip,pick}^{i}$ for $i=1,\cdots,N$.
-On the other hand, for placement, we did not make any assumptions about the orientation of the hole-platform structure.
-The heuristic we used for the place is as follows:
-$p_{place}=\bar{p}_{tip,place}-\bar{R}_{place}^{T}p_{et,place},$
-where $\bar{R}_{place}$ is the averaged rotation of $R_{place}^{i}$, and $\bar{p}_{tip,place}$ is the averaged position of the tip calculated similarly from the picking case.
-S-C Tips on applying Geometric Admittance Control
-Here, we present a few tips on the application of geometric admittance control. The GAC is implemented as a part of a ROS2 node, running with $200\mathrm{Hz}$.
-S-C1 ROS2 Implementation
-The Robot Node acts as the central controller, receiving and executing desired pose commands using a Geometric Admittance Control (GAC).
-Its main execution loop operates at $200\mathrm{Hz}$, ensuring smooth and compliant motion in response to external interactions.
-The node subscribes to the “action” topic, which contains the pose command and admittance gains command.
-Additionally, the controller subscribes to the “rebias” topic in order to preserve the accuracy of force readings.
-When a message on this topic is received, a callback raises a flag, triggering a re-biasing procedure for the FT sensor during the next control cycle.
-Within its main loop, the controller sends the intended pose and processed FT sensor data to the GAC function, which generates the compliant commands for the robot’s low-level motion controller.
-S-C2 Handling Force-Torque Sensor Data
-The first consideration is how to handle the Force-Torque (FT) sensor attached to the end-effector.
-It is always favorable to subtract the bias of the FT sensor after finishing each task, such as picking up the peg.
-We collect the FT sensor value for $2$ seconds and take a mean value to use as a bias.
-The other is to filter the FT sensor.
-Since the FT sensor is noisy, it is strongly recommended to implement a low-pass filter.
-However, having a low cut-off frequency can lead to serious phase lag, resulting in so-called “bouncing” behavior.
-We designed a simple 1st-order low-pass filter and implemented it using a state-space representation in discrete time.
-S-C3 Implementation Tips
-As mentioned in Section.
-IV, the detailed implementation method depends on the low-level servo controller type.
-Typical controllers provided by robot manufacturers are joint position control, joint velocity control, end-effector position control, and end-effector velocity control.
-If a torque controller is readily available, then directly implementing a geometric impedance controller (GIC) is recommended, but joint friction compensation needs to be handled appropriately.
-Given the desired dynamics (9), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
-$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
-$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (16)
-where $T_{s}$ is a sampling time ($5\mathrm{ms}$ in our case for GAC) and $\hat{(\cdot)}$ denotes a hat-map.
-The end-effector velocity control can be straightforwardly applied from (S-C3), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command.
-One caveat here is the $V^{b}$, which is body-frame end-effector velocity.
-The common velocity provided by the manufacturer is defined on the Cartesian (Spatial) frame.
-If end-effector velocity is only offered in the Cartesian frame $V\in\mathbb{R}^{6}$, then
-$V^{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}V,$
-where $R$ is a rotation matrix from the current configuration.
-However, sometimes, it is provided in the Cartesian frame for translational velocity and the body frame for rotational velocity.
-The users should verify the frames of the velocities by running the robot in one direction and logging its values.
-For the joint space-based approach, one can use
-$\begin{split}\dot{q}_{d}(k)\!&=\!\dot{q}(k)\!+\!T_{s}\cdot J_{b}(k)^{-1}M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),\\ q_{d}(k)\!&=\!q(k)+T_{s}\cdot\dot{q}_{d}(k),\end{split}$
-where $J_{b}$ denotes a body-frame Jacobian matrix.
-$\dot{q}_{d}(k)$ is provided to joint-velocity controller and $q_{d}(k)$ is provided to the joint-position controller.
-The Jacobian matrix also needs to be identified concerning the end-effector velocity, i.e., if the end-effector velocity is given in the Cartesian frame, then the Jacobian matrix $J$ is also provided in the Cartesian frame.
-In this case, the body-frame Jacobian matrix can be obtained via
-$J_{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}J$
-When the Jacobian matrix is not provided, an alternative is to use the robotics toolbox from Python or Matlab.
-First, use the DH parameters provided by the manufacturer (or URDF), then build the Jacobian matrix function.
-The function can be cast into a lambda function in Python.
-For Python, the Jacobian matrix function can also be wrapped into executable .so file using Matlab code builder and executed in the loop for fast computation (well below $1\mathrm{ms}$).
-S-D Teleoperation Details
-The architecture of our teleoperation system is explained to make it easier to reproduce and adapt our work.
-The system is constructed modularly, with the human-operator interface decoupled from the robot controller.
-This allows greater adaptation to different robotic platforms and control algorithms.
-The teleoperation function is implemented by two ROS nodes: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
-The teleoperation node acts as the principal interface between the human operator and the robot controller.
-A 6-DOF SpaceMouse is used to record input, which is then converted into the necessary end-effector pose.
-The main teleop function of this node operates at a frequency of $200\mathrm{Hz}$ inside a timer loop.
-It adjusts the desired end-effector pose incrementally during each cycle by reading velocity commands from the SpaceMouse.
-The SpaceMouse provides velocity commands $V_{sm}=[v_{sm}^{T},\omega_{sm}^{T}]^{T}\in\mathbb{R}^{6}$, where $v_{sm}\in\mathbb{R}^{3}$ is translational velocity and $\omega_{sm}\in\mathbb{R}^{3}$ is rotational velocity.
-The change in desired position, $\Delta p$, is computed by scaling $v_{sm}$ with a position offset:
-$\Delta p=v_{sm}\cdot\mathtt{pos\_offset}$
-The target position is then updated as:
-$p_{d}(k+1)=p_{d}(k)+\Delta p$
-The desired end-effector orientation, $R_{d}(k+1)$, is updated by composing the previous rotation with the incremental rotation matrix, $\Delta R$.
-The small, incremental angles, $\Delta\theta$, are computed by scaling $w_{sm}$ with rotation offset:
-$\Delta\theta=w_{sm}\cdot\mathtt{rot\_offset}$
-The $\Delta\theta$ is then converted to $\Delta R$ using Euler angle as $\Delta R=\mathtt{euler2rotmat(\Delta\theta)}$. The target orientation is then updated as:
-$R_{d}(k+1)=R_{d}(k)\cdot\Delta R$
-This calculated pose is then published as a command for the robot controller at a frequency of $200\mathrm{Hz}$.
-To create a dataset of the teleoperation commands, a separate timer logs the same message at $30\mathrm{Hz}$.
-This implements a teleoperation that operates on the robot base frame for translational movement and the end-effector frame for rotational movement.
-This teleoperation node was created with multiple operation modes to improve usability.
-For end-effector alignment with tilted platforms, the operator can switch the SpaceMouse’s control frame between the end-effector frame and the robot’s base frame.
-Additionally, a dedicated button allows users to modify the sensitivity offsets for position and rotation, enabling users to alternate between a translation-focused and an orientation-focused mode.
-The operator can globally scale motion sensitivity by cycling through low, medium, and high speed levels.
-The system also includes specific features, such as an “insertion mode” that operates in the end-effector frame for the translational movement to help with high-precision alignment tasks, as well as requests to set the robot’s end effector at “home” or a randomized pose.
-Furthermore, the operator can dynamically adjust the parameters of the Geometric Admittance Controller via dedicated keyboard inputs.
-There are four gain profiles: Free mode applies high stiffness gains $(1000)$ to all axes for precise motion in open space; Contact mode uses low z-axis gain $(300)$ with high gain elsewhere $(1500)$ to exert less force on surfaces; Insertion mode reverses this, with high z-axis gain $(1500)$ for precision while allowing compliance in other directions $(300)$; and Compliant mode sets all gains low $(300)$ to make the end-effector highly responsive to external forces.
-These gain profiles are published as action messages, together with poses and gripper states.
-S-E G-CompACT Training Details
-During the training of the G-CompACT, we added the noise to the known fixed pose of the hole and peg.
-As mentioned earlier, this is because the output of the Diff-EDFs $g_{{}_{EDF}}$ is biased.
-By adding the noise to the goal pose signal, we encourage the model to focus on the image of the wrist camera instead of the incorrect GCEV $e_{{}_{G}}$ signal.
-For our training, we added noise to every step of the demonstration dataset drawn from a uniform distribution.
-The noisy target poses $\tilde{g}_{ref}=(\tilde{p}_{ref},\tilde{R}_{ref})$are calculated as
-$\begin{split}\tilde{p}_{ref}&=p_{ref}+n_{p},\\ \tilde{R}_{ref}&=R_{ref}\cdot\mathtt{euler2rotmat}(n_{r}),\end{split}$
-where $n_{p}\sim\mathcal{U}(-0.02,0.02)\in\mathbb{R}^{3}$ and $n_{r}\sim\mathcal{U}(-8^{\circ},8^{\circ})\in\mathbb{R}^{3}$, which denotes $2\mathrm{cm}$ of translational and $8^{\circ}$ (degree) of rotational error.
-The impact of adding noise to the reference frame is further shown in Appendix.
--C.
-To encode the element on $SE(3)$ for learning, we used position & rotation vector representation for G-CompACT.
-However, in benchmark models of CompACT and ACTs, we utilized position & rot6d representation.
-This is because the rotation vector representation has a sign flip issue when the rotation angle is near $180^{\circ}$, but the rot6d representation does not have that problem.
-For G-CompACT, as the actions are represented with respect to the current end-effector frame, the transformation angles are relatively small and do not have a sign flip problem.
+Figure A2: Whole pipeline implemented with ROS2 is presented.
+As mentioned earlier, the full EquiContact pipeline is implemented with ROS2 framework. The whole implementation flow is presented in Fig. A2, and also summarized as in Algorithm 1.
+AIII Additional Experimental Results
+AIII-A Errors of Diff-EDF
+The RMSE error of the Diff-EDF on the training dataset is presented in Table A2. The RMSE of the rotational error is naively calculated from the Euler angles of the error rotation matrix.
+TABLE A2: RMSE error values of Diff-EDFs on the training dataset.
+The dimensions of translational errors in $x,y,z$ directions, given by $e_{T,x},e_{T,y},e_{T,z}$, are $\mathrm{mm}$ and the rotational errors in $x,y,z$ directions, given by $e_{R,x},e_{R,y},e_{R,z}$, are $\mathrm{deg}$.
+$e_{T,x}$ $e_{T,y}$ $e_{T,z}$ $e_{R,x}$ $e_{R,y}$ $e_{R,z}$
+pick 7.173 6.933 6.199 7.650 15.90 15.67
+place 13.75 8.241 5.999 3.806 5.560 5.660
+As noticed from the table, the translational error is significantly larger than the desired accuracy of precision of the PiH task $\sim 1\mathrm{mm}$.
+In addition, the rotational error of the picking task is significantly higher than that of the placing task.
+Therefore, we use the “upright peg” assumption for the full pipeline implementation.
+AIII-B Vision Encoder Design Study
+Here, we conduct a controlled comparison of vision encoder variants.
+To verify the design choices to meet the conditions of Assumption 1, we have trained $4$ models with the same training dataset for PiH tasks, which are listed below:
+  • Baseline ACT architecture that uses ResNet 18 and without language feature (RN18)
+  • ACT with pretrained CLIP-RN50 but is frozen (CLIP-RN50-frozen)
+  • ACT with CLIP-RN50, but $10\%$ of learning rate for vision backbone (CLIP-RN50-SB, SB stands for slow backbone training)
+  • ACT with CLIP-RN50, same learning rate for policy and vision backbone (proposed)
+We have tested our models in the in-distribution condition and with a $45^{\circ}$ transformation in the $y$ axis, i.e., the third case for extreme task transformations (Fig.
+3).
+The results are summarized in Table A3.
+TABLE A3: Results of vision encoder design study. OOD case here is a $45^{\circ}$ transformation in the $y$ axis.
+Backbone Learning Rate Learning Rate Success Rate
+($\eta_{policy}$) ($\eta_{vision}$) In-dist OOD
+RN18 1e051e-05 1e051e-05 10 / 10 6 / 10
+CLIP-RN50-frozen 1e051e-05 0 3 / 10 3 / 10
+CLIP-RN50-SB 1e051e-05 1e061e-06 10 / 10 0 / 10
+CLIP-RN50 (proposed) 1e051e-05 1e051e-05 10 / 10 10 / 10
+We first observe that the vision encoder without language guidance degrades under the OOD rotation (6/10), although background randomization during data collection partially mitigates background overfitting.
+In contrast, using a frozen CLIP-RN50 encoder yields low success even in-distribution (3/10), suggesting a significant domain mismatch between internet-scale pretraining and the short-range wrist-camera viewpoint in contact-rich manipulation.
+Interestingly, fine-tuning the CLIP-RN50 encoder with a very small learning rate achieves high in-distribution performance (10/10) but fails completely under the OOD rotation (0/10).
+We speculate that the visual representation is not sufficiently adapted: the encoder adjusts only locally to the training task configuration, without acquiring robustness to large geometric shifts, making the downstream policy brittle when viewpoint changes substantially.
+Finally, jointly fine-tuning the CLIP-RN50 encoder together with the policy (proposed) recovers both in-distribution and OOD performance (10/10), indicating that stronger encoder adaptation is critical for wrist-camera generalization under task transformations.